### v1

# Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting

Abstract
Test-time adaptation (TTA) seeks to tackle potential distribution shifts between training and testing data by adapting a given model w.r.t. any testing sample. This task is particularly important when the test environment changes frequently. Although some recent attempts have been made to handle this task, we still face two key challenges: 1) prior methods have to perform backpropagation for each test sample, resulting in unbearable optimization costs to many applications; 2) while existing TTA solutions can significantly improve the test performance on out-of-distribution data, they often suffer from severe performance degradation on in-distribution data after TTA (known as catastrophic forgetting). To this end, we have proposed an Efficient Anti-Forgetting Test-Time Adaptation (EATA) method which develops an active sample selection criterion to identify reliable and non-redundant samples for test-time entropy minimization. To alleviate forgetting, EATA introduces a Fisher regularizer estimated from test samples to constrain important model parameters from drastic changes. However, in EATA, the adopted entropy loss consistently assigns higher confidence to predictions even when the samples are underlying uncertain, leading to overconfident predictions that underestimate the data uncertainty. To tackle this, we further propose EATA with Calibration (EATA-C) to separately exploit the reducible model uncertainty and the inherent data uncertainty for calibrated TTA. Specifically, we compare the divergence between predictions from the full network and its sub-networks to measure the reducible model uncertainty, on which we propose a test-time uncertainty reduction strategy with divergence minimization loss to encourage consistent predictions instead of overconfident ones. To further re-calibrate predicting confidence on different samples, we utilize the disagreement among predicted labels as an indicator of the data uncertainty. Based on this, we devise a min-max entropy regularization to selectively increase and decrease predicting confidence for confidence re-calibration. Note that EATA-C and EATA are different on the adaptation objective, while EATA-C still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA. Extensive experiments on image classification and semantic segmentation verify the effectiveness of our proposed methods.
Index Terms:
Out-of-Distribution Generalization, Test-Time Adaptation, Calibration, Catastrophic, Forgetting, Robustness.
1 Introduction
Deep neural networks (DNNs) have achieved excellent performance in many challenging tasks, including image classification [1], video recognition [2, 3, 4, 5], and many other areas [6, 7, 8]. One prerequisite behind the success of DNNs is that the test samples are drawn from the same distribution as the training data, which, however, is often violated in many real-world applications. In practice, test samples may encounter natural variations or corruptions (also called distribution shift), such as changes in lighting resulting from weather changes and unexpected noises resulting from sensor degradation [9, 10]. Unfortunately, models are often very sensitive to such distribution shifts and suffer severe performance degradation.
Recently, several attempts [11, 12, 13, 14, 15, 16] have been proposed to handle the distribution shifts by online adapting a model at test time (called test-time adaptation). Test-time training (TTT) [11] first proposes this pipeline. Given a test sample, TTT first fine-tunes the model via rotation classification [17] and then makes a prediction using the updated model. Without the need of training an additional self-supervised head, Tent [12] and MEMO [14] further leverage the prediction entropy for test-time adaptation, in which the adaptation only involves test samples and a trained model. Although recent test-time adaptation methods are effective at handling test shifts, in real-world applications, they still suffer from the following limitations.
TABLE I: Characteristics of problem settings that adapt a trained model to a potentially shifted test domain. ‘Offline’ adaptation assumes access to the entire source or target dataset, while ‘Online’ adaptation can predict a single or batch of incoming test samples immediately.
Setting Source Data Target Data Training Loss Testing Loss Offline Online Source Acc. Prediction Uncertainty
Fine-tuning ${\bf x}^{t},y^{t}$ ${\mathcal{L}}({\bf x}^{t},y^{t})$ Not Considered Not Considered
Continual learning ${\bf x}^{t},y^{t}$ ${\mathcal{L}}({\bf x}^{t},y^{t})$ Maintained Not Considered
Unsupervised domain adaptation ${\bf x}^{s},y^{s}$ ${\bf x}^{t}$ ${\mathcal{L}}({\bf x}^{s},y^{s})+{\mathcal{L}}({\bf x}^{s},{\bf x}^{t})$ Maintained Not Considered
Test-time training ${\bf x}^{s},y^{s}$ ${\bf x}^{t}$ ${\mathcal{L}}({\bf x}^{s},y^{s})+{\mathcal{L}}({\bf x}^{s})$ ${\mathcal{L}}({\bf x}^{t})$ Not Considered Not Considered
Fully test-time adaptation (FTTA) ${\bf x}^{t}$ ${\mathcal{L}}({\bf x}^{t})$ Not Considered Not Considered
EATA (ours) ${\bf x}^{t}$ ${\mathcal{L}}({\bf x}^{t})$ Maintained Not Considered
EATA-C (ours) ${\bf x}^{t}$ ${\mathcal{L}}({\bf x}^{t})$ Maintained Calibrated
Latency Constraints. Since TTA adapts a given model during inference, the adaptation efficiency is paramount in scenarios where latency is a critical factor. Previous methods, such as Test-Time Training (TTT) [11] and MEMO [14], often require performing multiple backward propagations for each test sample. However, the computation-intensive nature of backward propagation renders these methods impractical in situations where low latency is non-negotiable or computational resources are limited.
Forgetting on In-Distribution Samples. Prior methods often focus on boosting the performance of a trained model on out-of-distribution (OOD) test samples, ignoring that the model after test-time adaptation suffers a severe performance degradation (named forgetting) on in-distribution (ID) test samples (see Figure 3). An eligible test-time adaptation approach should perform well on both OOD and ID test samples simultaneously, since test samples often come from both ID and OOD domains in real-world applications.
Over-Confident Predictions. Existing methods like Tent [12] and SAR [18] primarily focus on reducing the model’s prediction uncertainty through unsupervised prediction entropy minimization, which enhances the model’s predicting confidence, thereby learning a decision boundary positioned in low-density sample regions [19, 20]. However, a critical oversight of these methods is their neglect of the irreducible data uncertainty inherent in test samples. They encourage highly certain predictions for all test samples even in cases of underlying data uncertainty or even if the prediction is wrong, leading to overly certain (or overconfident) predictions. This phenomenon is particularly concerning in high-risk applications [21, 22], posing potential safety risks.
To address the efficiency and forgetting issue, we have proposed an Efficient Anti-forgetting Test-time Adaptation (EATA) method consisting of a sample-efficient optimization strategy and a weight regularizer. EATA excludes unreliable samples characterized by high entropy values and redundant samples that are highly similar throughout the adaptation. In this case, we can reduce the total number of backward updates of test data streaming (improving efficiency) and enhance the model performance on OOD samples. Furthermore, EATA devises an anti-forgetting regularizer to prevent the important weights of the model from changing a lot during the adaptation, where the weights’ importance is measured based on Fisher information [23] via a small set of test samples. With this regularization, the model can continually adapt to OOD test samples without performance degradation on ID test samples.
To mitigate the over-confident issue, it is essential to differentiate between various origins of uncertainty in TTA, i.e., reducible model uncertainty and inherent data uncertainty. Inspired by the human tendency to base predictions on available knowledge and problem difficulty, rather than defaulting to high certainty in unfamiliar situations, we propose to utilize the model uncertainty for calibrated uncertainty reduction and leverage the inherent data uncertainty to directly re-calibrate predicting confidence.
To this end, we further propose EATA with Calibration, namely EATA-C. To be specific, EATA-C estimates model uncertainty by measuring the divergence between predictions from the full network and its randomly sampled sub-networks. This allows for the reduction of model uncertainty by minimizing the divergence, promoting consistent rather than overconfident predictions during testing. Additionally, we develop a simple but effective indicator of data uncertainty based on the disagreement among predicted labels, which suggests whether a sample nears decision boundaries. Then, we devise an auxiliary min-max entropy regularizer to selectively lower confidence for inherently uncertain samples through an entropy maximization loss, while simultaneously boosting confidence for certain samples via an entropy minimization loss. EATA-C differs from EATA on the adaptation objective, while it still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA. We summarize our main contributions as follows.
  • We propose an Efficient Anti-forgetting Test-time Adaptation (EATA) method. Specifically, we develop an active sample identification scheme to filter out non-reliable and redundant test samples from model adaptation, thereby improving adaptation efficiency by reducing backward propagation. Moreover, we extend the label-dependent Fisher regularizer to test samples with pseudo label generation, which prevents drastic changes in important model weights and thus alleviates the issue of model forgetting on in-distribution test samples.
  • We further propose EATA with Calibration (EATA-C), which decouples the model and data uncertainty to achieve confidence-calibrated TTA. Specifically, EATA-C estimates model uncertainty using the divergence between full network and sub-network predictions, incorporating a consistency loss to reduce this uncertainty without leading to overconfidence. For data uncertainty, EATA-C utilizes label prediction disagreements, employing min-max entropy regularization to selectively adjust prediction confidence for effective re-calibration.
  • We demonstrate that our proposed EATA method improves both the performance and efficiency of test-time adaptation and also alleviates the long-neglected catastrophic forgetting issue, and EATA-C further achieves better performance and calibration without additional backward computations compared with EATA.
Figure 1: An illustration of our proposed Efficient Anti-forgetting Test-time Adaptation with Calibration (EATA-C) method. During the test-time adaptation process, we update only the affine parameters of normalization layers in $f_{\Theta}$ and keep all other parameters frozen. Given a batch of incoming test samples ${\mathcal{X}}\small{=}\{{\bf x}_{b}\}_{b=1}^{B}$, we select the reliable and non-redundant ones ${\mathcal{X}}_{s}$ with an active sample selection criterion to conduct model update, thereby enhancing adaptation efficiency. These samples are then used for calculating the proposed unidirectional consistency loss to minimize the model uncertainty. Additionally, we devise a min-max entropy regularizer for confidence re-calibration based on the data uncertainty of each sample. Lastly, we introduce an anti-forgetting regularizer which prevents the important model parameters in $\Theta$ from changing too much.
2 Related Work
We divide the discussion on related works based on the different adaptation settings summarized in Table I and further review existing methods for model’s uncertainty calibration.
Test-Time Adaptation (TTA) aims to improve model accuracy on OOD test data through model adaptation with test samples, which has been successfully applied to various tasks, such as image classification [11, 12], image super-resolution [24], image captioning [25], object detection [26], video recognition [27, 28], etc. Existing test-time training methods, e.g., TTT [11], TTT++ [13], TTT-MAE [29], and MT3 [30], jointly train a source model via both supervised and self-supervised objectives, and then adapt the model via self-supervised objective at test time. This pipeline, however, necessitates both self-supervised head and test data in adaptation, while training such self-supervised head can be computation-consuming [29]. To address this, some methods have been proposed to adapt a model with only test data, including batchnorm statistics adaptation [31, 32, 33], prediction consistency maximization over different augmentations [34], and classifier adjustment [35]. Specifically, Tent [12] updates the model to minimize the entropy of predictions at test time. MEMO [14] further augments test samples for marginal entropy minimization to enhance robustness. Our work also alleviates the dependency on self-supervision heads and seeks to address the key limitations of prior works (i.e., efficiency hurdle, catastrophic forgetting, and overconfidence) to make TTA more practical in real-world applications.
Continual Learning (CL) aims to help the model remember the essential concepts that have been learned previously, alleviating the catastrophic forgetting issue when learning a new task [23, 36, 37, 38, 39, 40]. In our work, we share the same motivation as CL and point out that test-time adaptation also suffers catastrophic forgetting (i.e., performance degradation on ID test samples), which makes TTA approaches unstable to deploy. To conquer this, we propose a simple yet effective solution to maintain the model performance on ID test samples (by only using test data) and meanwhile improve the performance on OOD test samples.
Unsupervised Domain Adaptation (UDA). Conventional UDA tackles distribution shifts by jointly optimizing a source model on both labeled source data and unlabeled target data, such as devising a domain discriminator to learn domain-invariant features [41, 42, 43, 44]. To avoid access to source data, recently CPGA [45] generates feature prototypes for each category with pseudo-labeling. SHOT [46] learns a target-specific feature extractor by information maximization for representations alignment. Nevertheless, such methods optimize offline via multiple epochs and losses. In contrast, our method adapts in an online manner and selectively performs once backward propagation for one given target sample, which is more efficient during inference.
Uncertainty Calibration. A calibrated model refers to whose predicting confidence reflects the true likelihood of correctness. Post-training processing methods [47, 48, 49] re-calibrate a trained model by leveraging a labeled dataset within the target domain to estimate calibration error. In contrast, regularization-based methods [50, 51, 52, 53] introduce auxiliary objectives to improve calibration at the training phase. Recently, SB-ECE [54] proposes a differentiable estimation of calibration error as regularization to be jointly minimized. ESD [55] further reformulates the calibration objective in a class-wise manner to enhance calibration performance. Nevertheless, these methods necessitate labeled data from the source or target domain, which limits their applicability. Unlike these methods, we seek to improve calibration with only access to unlabeled test data in an online manner in TTA context.
3 Problem Formulation
Without loss of generality, let $P\left({\bf x}\right)$ be the distribution of training data $\{{\bf x}_{i}\}_{i=1}^{N}$ (namely ${\bf x}_{i}\sim P\left({\bf x}\right)$) and $f_{\Theta^{o}}({\bf x})$ be a base model trained on labeled training data $\{({\bf x}_{i},y_{i})\}_{i=1}^{N}$, where $\Theta^{o}$ denotes the model parameters. Due to the training process, the model $f_{\Theta^{o}}({\bf x})$ tends to fit (or overfit) the training data. During the inference state, the model shall perform well for the in-distribution test data, namely ${\bf x}\sim P\left({\bf x}\right)$. However, in practice, due to possible distribution shifts between training and test data, we may encounter many out-of-distribution test samples, namely ${\bf x}\sim Q\left({\bf x}\right)$, where $Q\left({\bf x}\right)\neq P\left({\bf x}\right)$. In this case, the prediction would be very unreliable and the performance is also very poor.
Test-time adaptation (TTA) [12, 14] aims at boosting the out-of-distribution prediction performance by doing model adaptation on test data only. Specifically, given a set of test samples $\{{\bf x}_{j}\}_{j=1}^{M}$, where ${\bf x}_{j}\sim Q\left({\bf x}\right)$ and $Q\left({\bf x}\right)\neq P\left({\bf x}\right)$, one needs to adapt $f_{\Theta}({\bf x})$ to improve the prediction performance on test data in any cases. To achieve this, existing methods often seek to update the model by minimizing some unsupervised objective defined on test samples:
$\min_{\tilde{\Theta}}{\mathcal{L}}({\bf x};\Theta),~{}{\bf x}\sim Q\left({\bf x}\right),$ (1)
where $\tilde{\Theta}\subseteq\Theta$ denotes the free model parameters that should be updated. In general, the test-time learning objective ${\mathcal{L}}(\cdot)$ can be formulated as an entropy minimization problem [12] or prediction consistency maximization over data augmentations [14], etc..
For existing TTA methods like TTT [11] and MEMO [14]), during the test-time adaptation, we shall need to compute one round or even multiple rounds of backward computation for each sample, which is very time-consuming and also not favorable for latency-sensitive applications. Moreover, most methods assume that all the test samples are drawn from out-of-distribution (OOD). However, in practice, the test samples may come from both in-distribution (ID) and OOD. Simply optimizing the model on OOD test samples may lead to severe performance degradation on ID test samples. We empirically validate the existence of this issue in Figure 3, where the updated model has a consistently lower accuracy on ID test samples than the original model.
Moreover, existing entropy-based test-time adaptation methods like Tent [12] and SAR [18] consistently encourage producing one-hot predictions with high confidence. However, in practice, incoming test samples can be severely corrupted [9], inheriting irreducible data uncertainty. Considering the data uncertainty, these test samples ideally should be predicted with relatively low confidence. Nevertheless, the data uncertainty is often overlooked by methods based on entropy minimization. Consequently, the adapted model often produces highly confident predictions (called overconfidence), even when the test samples are uncertain. Such misleading predictions raise potential safety concerns for real-world application scenarios. We empirically demonstrate the issue of overconfidence in Figure 7(a) and Table II.
4 Uncertainty-Calibrated Efficient Anti-forgetting Test-time Adaptation
In this section, we first propose an Efficient Anti-forgetting Test-time Adaptation (EATA) method, which aims to improve the efficiency of test-time adaptation (TTA) and tackle the catastrophic forgetting issue brought by existing TTA strategies simultaneously. EATA consists of two strategies. 1) Sample-efficient entropy minimization (c.f. Section 4.1) aims to conduct efficient adaptation relying on an active sample selection strategy. Here, the sample selection process is to choose only active samples for backward propagation and therefore improve the overall TTA efficiency (i.e., less gradient backward propagation). To this end, we devise an active sample selection score, denoted by $S({\bf x})$, to detect those reliable and non-redundant test samples from the test set for TTA. 2) Anti-forgetting weight regularization (c.f. Section 4.2) seeks to alleviate knowledge forgetting by enforcing that the parameters, important for the ID domain, do not change too much in TTA. In this way, the catastrophic forgetting issue can be significantly alleviated. We illustrate EATA in Figure A in Supplementary.
To further address the overconfidence issue, we propose an Efficient Anti-forgetting Test-time Adaptation with Calibration (EATA-C) method. As shown in Figure 1, we introduce a new consistency-based test-time learning objective for model uncertainty reduction (c.f. Section 4.3), follow up a min-max entropy regularizer to re-calibrate the prediction uncertainty according to the inherent data uncertainty (c.f. Section 4.4).
4.1 Sample Efficient Entropy Minimization
For efficient test-time adaptation, we propose an active sample identification strategy to select samples for backward propagation. Specifically, we design an active sample selection score for each sample, denoted by $S({\bf x})$, based on two criteria: 1) samples should be reliable for test-time adaptation, and 2) the samples involved in optimization should be non-redundant. By setting $S({\bf x})\small{=}0$ for non-active samples, namely the unreliable and redundant samples, we can reduce unnecessary backward computation during test-time adaptation, thereby improving the prediction efficiency.
Relying on the sample score $S({\bf x})$, following [12, 14], we use entropy loss for model adaptations. Then, the sample-efficient entropy minimization is to minimize the following objective:
$\min_{\tilde{\Theta}}S({\bf x})E({\bf x};\Theta)\small{=-}S({\bf x})\sum_{y\in{\mathcal{C}}}f_{\Theta}(y|{\bf x})\text{log}f_{\Theta}(y|{\bf x}),$ (2)
where ${\mathcal{C}}$ is the model output space. Here, the entropy loss $E(\cdot)$ is calculated over a batch of samples each time (similar to Tent [12]) to avoid a trivial solution, i.e., assigning all probability to the most probable class. For efficient adaptation, we update $\tilde{\Theta}\small{\subseteq}\Theta$ with the affine parameters of all normalization layers.
Reliable Sample Identification. Our intuition is that different test samples produce various effects in adaptation. To verify this, we conduct a preliminary study, where we select different proportions of samples (the samples are pre-sorted according to their entropy values $E({\bf x};\Theta)$) for adaptation, and the resulting model is evaluated on all test samples. From Figure 2, we find that: 1) adaptation on low-entropy samples makes more contribution than high-entropy ones, and 2) adaptation on test samples with very high entropy may hurt performance. The possible reason is that predictions of high-entropy samples are uncertain, so their gradients produced by entropy loss may be biased and unreliable. Following this, we name these low-entropy samples as reliable samples. Based on the above observation, we propose an entropy-based weighting scheme to identify reliable samples and emphasize their contributions during adaptation. Formally, the entropy-based weight is given by:
$S^{ent}({\bf x})=\frac{1}{\exp\left[E({\bf x};\Theta)-E_{0}\right]}\cdot\mathbb{I}_{\{E({\bf x};\Theta)<E_{0}\}}({\bf x}),$ (3)
where $\mathbb{I}_{\{\cdot\}}(\cdot)$ is an indicator function, $E({\bf x};\Theta)$ is the predicted entropy regarding sample ${\bf x}$, and $E_{0}$ is a pre-defined threshold. The above weighting function excludes high-entropy samples from adaptation and assigns higher weights to test samples with lower prediction uncertainties, allowing them to contribute more to model updates. Note that evaluating $S^{ent}({\bf x})$ does not involve any gradient back-propagation.
Figure 2: Effect of different test samples in test-time entropy minimization [12]. We adapt a model on partial samples (top $p$samples with the highest or lowest entropy values), and then evaluate the adapted model on all test samples. Results are obtained on ImageNet-C (Gaussian noise, level 3) and ResNet-50 (base accuracy is 27.6%). Introducing more samples with high entropy values into adaptation will hurt the adaptation performance.
Non-redundant Sample Identification. Although Eqn. (3) helps to exclude partial unreliable samples, the remaining test samples may still have redundancy. For example, given two test samples that are mutually similar and both have a lower prediction entropy than $E_{0}$, we still need to perform gradient back-propagation for each of them according to Eqn. (3). However, this is unnecessary as these two samples produce similar gradients for model adaptation.
To further improve efficiency, we propose to exploit the samples that can produce different gradients for model adaptation. Recall that the entropy loss only relies on final model outputs (i.e., classification logits), we further filter samples by ensuring the remaining samples have diverse model outputs. To this end, one straightforward method is to save the model outputs of all previously seen samples, and then compute the similarity between the outputs of incoming test samples and all saved model outputs for filtering. However, this method is computationally expensive at test time and memory-consuming with the increase of test samples.
To address this, we exploit an exponential moving average technique to track the average model outputs of all seen test samples used for model adaptation. To be specific, given a set of model outputs of test samples, the moving average vector is updated recursively:
${\bf m}^{t}=\left\{\begin{array}[]{ll}\bar{{\bf y}}^{1},&\text{ if }t=1\\ \alpha\bar{{\bf y}}^{t}+(1-\alpha){\bf m}^{t-1},&\text{ if }t>1\end{array},\right.$ (6)
where $\bar{{\bf y}}^{t}=\frac{1}{n}\sum_{k=1}^{n}\hat{{\bf y}}_{k}^{t}$ is the average model prediction of a mini-batch of $n$ test samples at the iteration $t$, and $\alpha\in[0,1]$. Following that, given a new test sample ${\bf x}$ received at iteration $t>1$, we compute the cosine similarity between its prediction $f_{\Theta}({\bf x})$ and the moving average ${\bf m}^{t-1}$ (i.e., $cos(f_{\Theta}({\bf x}),{\bf m}^{t-1})$), which is then used to determine the diversity-based weight:
$S^{div}\left({\bf x}\right)=\mathbb{I}_{\{cos(f_{\Theta}({\bf x}),{\bf m}^{t-1})<\epsilon\}}({\bf x}),$ (7)
where $\epsilon$ is a pre-defined threshold for cosine similarities. The overall sample-adaptive weight is then given by:
$S\left({\bf x}\right)=S^{ent}\left({\bf x}\right)\cdot S^{div}\left({\bf x}\right),$ (8)
which combines both entropy-based (as in Eqn. 3) and diversity-based terms (as in Eqn. 7). Since we only perform gradient back-propagation for test samples with $S({\bf x})>0$, the algorithm efficiency is further improved.
Remark. Given $M$ test samples ${\mathcal{D}}_{test}=\{{\bf x}_{j}\}_{j=1}^{M}$, the total number of reduced backward computations is given by $\mathbb{E}_{{\bf x}\sim{\mathcal{D}}_{test}}[\mathbb{I}_{\{S({\bf x})=0\}}({\bf x})]$, which is jointly determined by test data ${\mathcal{D}}_{test}$, entropy threshold $E_{0}$, and cosine similarity threshold $\epsilon$.
4.2 Anti-Forgetting with Fisher Regularization
In this section, we propose a new weighted Fisher regularizer (called anti-forgetting regularizer) to alleviate the catastrophic forgetting issue caused by test-time adaptation, i.e., the performance of a test-time adapted model may significantly degrade on in-distribution (ID) test samples. We achieve this through weight regularization, which only affects the loss function and does not incur any additional computational overhead for model adaptation. To be specific, we apply an importance-aware regularizer ${\mathcal{R}}$ to prevent model parameters, important for the in-distribution domain, from changing too much during the test-time adaptation process [23]:
${\mathcal{R}}(\tilde{\Theta},\tilde{\Theta}^{o})=\sum_{\theta_{i}\in\tilde{\Theta}}\omega(\theta_{i})(\theta_{i}-\theta_{i}^{o})^{2},$ (9)
where $\tilde{\Theta}$ are parameters used for model update and $\tilde{\Theta}^{o}$ are the corresponding parameters of the original model. $\omega(\theta_{i})$ denotes the importance of $\theta_{i}$ and we measure it via the diagonal Fisher information matrix as in elastic weight consolidation [23]. Here, the calculation of Fisher information $\omega(\theta_{i})$ is non-trivial since we are inaccessible to any labeled training data. For the convenience of presentation, we leave the details of calculating $\omega(\theta_{i})$ in the next subsection.
After introducing the anti-forgetting regularizer, the final optimization formula for EATA is formulated as:
$\min_{\tilde{\Theta}}S({\bf x})E({\bf x};\Theta)+\beta{\mathcal{R}}(\tilde{\Theta},\tilde{\Theta}^{o}),$ (10)
where $\beta$ is a trade-off parameter, $S({\bf x})$ and $E({\bf x};\Theta)$ are defined in Eqn. (2).
Measurement of Weight Importance $\omega(\theta_{i})$. The calculation of Fisher information typically involves a set of labeled ID training samples. However, in our problem setting, we are inaccessible to training data and the test samples are only unlabeled, which makes it non-trivial to measure the weight importance. To conquer this, we first collect a small set of unlabeled ID test samples $\{{\bf x}_{q}\}_{q=1}^{Q}$, and then use the original trained model $f_{\Theta}(\cdot)$ to predict all these samples for obtaining the corresponding hard pseudo-label $\hat{y}_{q}$. Following that, we construct a pseudo-labeled ID test set ${\mathcal{D}}_{F}=\{{\bf x}_{q},\hat{y}_{q}\}_{q=1}^{Q}$, based on which we calculate the fisher importance of model weights by:
$\omega(\theta_{i})=\frac{1}{Q}\sum_{{\bf x}_{q}\in{\mathcal{D}}_{F}}\big{(}\frac{\partial}{\partial\theta_{i}^{o}}{\mathcal{L}}_{CE}(f_{\Theta^{o}}({\bf x}_{q}),\hat{y}_{q})\big{)}^{2},$ (11)
where ${\mathcal{L}}_{CE}$ is the cross-entropy loss. Here, we only need to calculate $\omega(\theta_{i})$ once before performing test-time adaptation. Once calculated, we keep $\omega(\theta_{i})$ fixed and apply it to any types of distribution shifts. Moreover, the unlabeled ID test samples are collected based on out-of-distribution detection techniques [56, 57], which are easy to implement. Note that there is no need to collect too many ID test samples for calculating $\omega(\theta_{i})$, e.g., 500 samples are enough for the ImageNet-C dataset. More empirical studies regarding this can be found in Figure 6.
4.3 Consistency-Based Uncertainty Minimization
As mentioned in Section 4.1, EATA enhances model performance by minimizing prediction entropy. This strategy aims to reduce uncertainty in predictions and learn a decision boundary in the low-density regions of the test samples [58, 59]. However, a persistent challenge with entropy minimization is its tendency to yield overly certain predictions, even in cases where the test samples themselves are inherently uncertain, such as severely corrupted images [9]. Thus, EATA may still result in overconfident predictions that fail to accurately account for the inherent data uncertainty. To address this, we further propose EATA with Calibration (EATA-C), shall be depicted in Sections 4.3 and 4.4.
In EATA-C, we first propose a consistency-based loss to quantify and optimize the model uncertainty. Our method is inspired by MC Dropout [60], which has shown promising performance in estimating the model uncertainty through the divergence of multiple dropout-enabled predictions. In our context, considering adaptation efficiency, we define the model uncertainty as the KL divergence [61] between the full network prediction and randomly sampled sub-network prediction. Here, we use only two predictions and the former is indispensable since we select it as the final prediction. Thus, at test time, we minimize divergence to encourage consistent predictions instead of confidence-related entropy loss, thereby alleviating the overconfidence issue.
Consistency Loss. Formally, let $\hat{{\bf y}}=f_{\Theta}({\bf x})$ be the prediction of the full network w.r.t. sample ${\bf x}$, and $\hat{{\bf y}}_{sub}=f_{\Theta_{sub}}({\bf x})$ be that of the sub-network. The consistency loss is defined as follows:
${\mathcal{L}}_{c}({\bf x})$ $=D_{KL}(\hat{{\bf y}}_{sub},\hat{{\bf y}}_{fuse}),$ (12)
$\hat{{\bf y}}_{fuse}=($ $\hat{{\bf y}}+(1-p)\cdot\hat{{\bf y}}_{sub})/(2-p),$ (13)
where $D_{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence [61] and $p$ is a constant for balancing. Here, we calculate the divergence between $\hat{{\bf y}}$ and $\hat{{\bf y}}_{fuse}$ (rather than $\hat{{\bf y}}_{sub}$), since encouraging the sub-network to achieve the same performance as the full one is relatively hard. Thus, inspired by label smoothing [62], we softly fuse $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ in Eqn. (13) for divergence optimization. Note that, during optimization, we conduct a unidirectional alignment from the sub-network to the full network, as the full network typically exhibits stronger generalization capabilities. To this end, we detach the gradient from $\hat{{\bf y}}$ and concentrate the optimization solely on $\hat{{\bf y}}_{sub}$. This strategy is designed to facilitate knowledge transfer from the full network to its sub-network, thereby enhancing the sub-network’s performance while simultaneously reducing the model uncertainty of the full network.
Remark on Efficiency. Although consistency loss requires two forward passes from both the full network and the sub-network for each sample, the full network’s forward pass is gradient-free without back-propagation and the sub-network’s forward/backward pass is less computationally intensive. Moreover, we only perform sub-network’s forward/backward pass on the selected reliable and non-redundant samples as outlined in Algorithm 1. As a result, the use of consistency loss remains efficient.
Algorithm 1 The pipeline of proposed EATA and EATA-C.
0:  Test samples ${\mathcal{D}}_{test}\small{=}\{{\bf x}_{j}\}_{j=1}^{M}$, the trained model $f_{\Theta}(\cdot)$, ID samples ${\mathcal{D}}_{F}\small{=}\{{\bf x}_{q}\}_{q=1}^{Q}$, batch size $B$.
1:  for a batch ${\mathcal{X}}\small{=}\{{\bf x}_{b}\}_{b=1}^{B}$ in ${\mathcal{D}}_{test}$ do
2:     Calculate predictions $\hat{y}$ for all ${\bf x}\in{\mathcal{X}}$ via $f_{\Theta}(\cdot)$.
3:     For EATA:
4:        Calculate sample selection score $S({\bf x})$ via Eqn. (8).
5:        Update model ($\tilde{\Theta}\subseteq\Theta$) with Eqn. (10).
6:     For EATA-C:
7:        Select reliable and distinct samples ${\mathcal{X}}_{s}$ via Eqn. (17).
8:        Sample $f_{\Theta_{sub}}(\cdot)$ from $f_{\Theta}(\cdot)$ via stochastic depth [63].
9:        Calculate predictions $\hat{y}_{sub}$ for ${\bf x}\in{\mathcal{X}}_{s}$ via $f_{\Theta_{sub}}(\cdot)$.
10:        Compute consistency loss ${\mathcal{L}}_{c}({\bf x})$ based on Eqn. (12)
11:        Calibrate confidence with entropy loss via Eqn. (14)
12:        Update model ($\tilde{\Theta}\subseteq\Theta$) with Eqn. (16).
13:  end for
13:  The predictions $\{\hat{y}\}_{j=1}^{M}$ for all ${\bf x}\in{\mathcal{D}}_{test}$.
4.4 Calibrated Min-Max Entropy Regularization
In this section, we re-calibrate the model’s prediction uncertainty in a manner that is sensitive to individual samples. This process involves categorizing samples into two distinct groups—‘certain’ and ‘uncertain’—based on the aforementioned prediction consistency. This identification is achieved by comparing the predicted labels between the full network and a sub-network. Samples exhibiting mismatched predictions are deemed ‘uncertain’, indicating either high intrinsic data uncertainty or a high level of complexity that challenges the model’s recognition capability. For these uncertain samples, we aim to lower their predictive confidence. This is achieved by maximizing the prediction entropy, effectively acknowledging the model’s lack of confidence in these cases. Conversely, for samples where predictions are consistent, labeled as ‘certain’, we conduct the opposite strategy, i.e., boosting prediction confidence through entropy minimization. Formally, this min-max entropy regularization optimization problem is defined by:
$\min_{\tilde{\Theta}_{sub}}C({\bf x})E({\bf x};\Theta_{sub}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (14)
$C({\bf x})=\begin{cases}1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})=\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ -1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})\neq\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ \end{cases}$ (15)
where $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ denote the prediction of the full and sub-network respectively, $\Theta_{sub}$ denotes the parameters of the sub-network and $\tilde{\Theta}_{sub}\subset\Theta_{sub}$ denotes parameters involved in model adaptation. Note that we only update the affine parameters of the sub-network considering efficiency as mentioned in Section 4.3.
Overall Objective. The methods proposed in Sections 4.3 and 4.4 are devised to address the overconfident issue in TTA, but still suffer from catastrophic forgetting when the important model weights for the in-distribution domain are modified significantly during adaptation. Therefore, we jointly optimize the model with the anti-forgetting regularizer and further reduce the required backward computations with the active sample selection criterion proposed in EATA. Then, the overall objective of EATA-C can be formulated as:
$\min_{\tilde{\Theta}}S_{c}({\bf x})\Big{(}{\mathcal{L}}_{c}({\bf x})+\alpha C({\bf x})E({\bf x};\Theta_{sub})\Big{)}+\beta{\mathcal{R}}(\tilde{\Theta},\tilde{\Theta}^{o}).$ (16)
where $\alpha$ and $\beta$ are balance factors, ${\mathcal{R}}(\tilde{\Theta},\tilde{\Theta}^{o})$ is the fisher regularizer defined in Eqn. (9), and $S_{c}({\bf x})$ is the joint indicator function in Eqn. (3) and Eqn. (7) to select reliable and non-redundant samples. To be specific, $S_{c}({\bf x})$ is defined by:
$S_{c}({\bf x})=S^{div}({\bf x})\cdot\mathbb{I}_{\{E({\bf x};\Theta)<E_{0}\}}({\bf x}).$ (17)
We summarize the overall pipeline of our proposed EATA-C and EATA in Algorithm 1.
5 Experiments
TABLE II: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$). “BN” and “LN” denote batch and layer normalization, respectively. The bold number indicates the best result and the underlined number indicates the second best result. All results are evaluated under the lifelong adaptation scenario except for Tent [12] and MEMO [14], which suffer severely from error accumulation. We use * and ${}^{\dagger}$ to denote episodic and single-domain adaptation, respectively.
Noise Blur Weather Digital Average
Model Method Metric Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel JPEG Avg. #Forwards #Backards
R-50 (BN) Source Acc. 1.8 3.0 1.7 18.2 10.1 13.4 20.8 14.0 22.1 21.9 58.7 5.3 17.6 22.1 37.5 17.88 50,000 0
ECE 16.6 16.1 15.9 1.8 10.7 10.7 14.7 25.3 12.9 16.7 2.3 6.7 22.9 10.5 6.0 12.64
BN Adapt Acc. 15.8 16.7 15.3 18.7 19.3 29.8 41.7 35.8 35.0 50.5 65.9 18.1 49.3 51.7 42.0 33.70 50,000 0
ECE 1.1 0.8 1.0 3.0 1.3 0.8 3.4 1.1 1.0 5.4 1.4 7.6 4.3 3.8 4.8 2.72
${}^{\dagger}$ Acc. 28.0 30.1 28.1 29.9 29.5 42.2 49.7 46.2 41.5 57.7 67.1 30.0 55.7 58.3 52.5 43.11 50,000 50,000
ECE 11.7 11.2 11.1 12.6 12.3 7.7 5.4 6.5 8.8 3.4 2.9 21.9 3.5 3.7 4.0 8.46
MEMO* Acc. 6.8 8.5 7.5 20.5 13.4 19.8 25.8 22.1 27.7 27.6 60.9 11.3 24.4 32.2 37.9 23.09 50,000$\times$65 50,000$\times$64
ECE 24.1 24.2 22.9 5.3 19.3 14.8 23.4 30.4 18.7 24.6 7.2 14.9 29.4 19.3 13.6 19.47
CoTTA Acc. 19.9 31.8 35.3 30.4 34.4 40.2 43.2 39.3 38.6 47.7 51.8 36.0 43.5 46.7 43.1 38.80 152,315 50,000
ECE 3.6 17.4 21.5 27.8 30.4 31.2 31.6 33.2 34.2 32.5 33.1 36.0 34.2 36.1 35.9 29.25
SAR Acc. 29.6 38.4 37.8 31.5 32.8 41.4 48.6 42.9 40.2 53.3 63.7 37.7 53.0 56.3 52.3 43.96 85,964 68,145
ECE 3.7 7.4 9.1 14.4 15.1 12.6 9.6 11.7 13.3 8.4 7.0 16.8 8.5 8.4 8.5 10.29
EATA (Ours) Acc. 35.6 38.7 37.5 35.9 36.1 47.6 53.1 50.6 45.6 59.5 67.1 45.2 57.6 59.8 55.4 48.36
ECE 10.5 13.4 14.7 18.5 18.5 14.7 12.8 14.1 16.1 11.6 10.4 18.6 13.3 13.1 13.9 14.28 50,000 29,721
EATA-C (Ours) Acc. 37.2 40.9 39.9 36.6 37.1 48.5 52.7 51.9 47.2 60.4 67.0 49.0 57.6 60.1 56.1 49.48
ECE 7.1 6.7 7.2 10.1 9.5 5.9 4.9 4.8 5.8 3.2 3.2 6.0 4.0 3.9 3.9 5.74 83,312 33,312
ViT (LN) Source Acc. 12.9 17.6 11.7 34.4 27.7 43.7 36.2 43.4 45.4 52.8 73.3 45.5 37.9 54.7 60.2 39.84 50,000 0
ECE 14.2 11.3 15.2 2.2 8.2 6.0 10.7 7.5 8.3 5.0 2.1 2.3 12.2 4.1 2.9 7.48
${}^{\dagger}$ Acc. 33.4 42.1 41.4 48.8 45.2 54.9 48.2 55.6 55.1 64.4 75.2 62.5 51.6 65.5 65.0 53.92 50,000 50,000
ECE 24.1 10.1 11.7 8.6 10.4 7.5 12.4 8.0 8.0 4.8 2.4 5.2 10.0 4.3 4.1 8.78
MEMO* Acc. 32.2 35.1 32.6 37.5 28.5 43.3 40.1 45.2 47.0 53.9 73.3 53.3 39.6 59.5 62.8 45.60 50,000$\times$65 50,000$\times$64
ECE 33.0 32.5 33.2 36.7 45.7 41.0 47.2 39.8 38.8 33.7 20.0 26.7 48.5 30.4 27.2 35.63
CoTTA Acc. 45.6 58.8 58.4 40.1 50.9 50.4 44.4 46.1 51.4 52.2 57.7 36.5 53.9 55.8 55.4 50.51 50,000$\times$3 50,000
ECE 9.0 15.8 21.0 29.3 28.3 29.1 34.5 29.8 29.6 30.0 29.2 36.3 29.0 31.2 31.8 27.59
SAR Acc. 43.1 50.8 52.9 50.4 51.0 57.5 53.2 58.6 61.2 66.0 76.1 61.2 54.7 67.8 67.7 58.15 91,605 82,277
ECE 8.0 8.4 9.5 9.1 11.4 9.6 13.0 9.9 9.6 7.8 4.5 10.0 13.8 7.5 7.5 9.30
EATA (Ours) Acc. 50.5 55.6 56.0 54.9 56.3 61.1 59.8 64.3 64.0 70.1 77.4 65.5 63.1 70.4 69.5 62.57
ECE 10.6 13.9 15.7 16.4 17.3 15.9 17.5 14.8 15.7 12.8 9.0 15.4 17.5 13.3 13.9 14.63 50,000 36,688
EATA-C (Ours) Acc. 56.8 60.2 59.8 58.0 60.9 65.2 65.5 69.7 67.8 74.0 78.9 66.7 70.3 74.1 71.8 66.65
ECE 5.2 5.1 5.9 7.1 6.2 5.6 5.9 5.0 5.0 4.1 3.3 5.4 4.7 4.2 4.7 5.14 83,184 33,184
TABLE III: Comparison on ImageNet-R. Results are evaluated in the single-domain adaptation scenario. We use * to denote episodic adaptation.
Model Acc. (%) ECE (%) #Forwards #Backwards
ResNet-50(BN) 38.0 17.7 30,000 0
[…]
ViT(LN) 52.5 5.0 30,000 0
[…]
Datasets and Models. We conduct experiments on three benchmarks datasets for OOD generalization, i.e., ImageNet-C [9] (contains corrupted images in 15 types of 4 main categories and each type has 5 severity levels) and ImageNet-R [65] for image classification, and ACDC [66] for semantic segmentation. We use ResNet-50 (R-50) [1] and VitBase (ViT) [67] for ImageNet experiments, and Segformer-B5 [68] for ACDC [66] experiments. The models are trained on ImageNet or CityScapes [69] training set with stochastic depth regularization [63] and then tested on clean or the above OOD test sets.
Compared Methods. We compare with the following state-of-the-art methods. BN adaptation [32] updates batch normalization statistics on test samples. Tent [12] minimizes the entropy of test samples during testing. MEMO [14] maximizes the prediction consistency of different augmented copies regarding a given test sample. CoTTA [64] minimizes the cross entropy between the predictions of the student network and its mean teacher at test time. SAR [18] selects reliable samples to perform test time sharpness-aware entropy minimization. We denote EATA without weight regularization in Eqn. (9) as efficient test-time adaptation (ETA). More ablative methods can be found in Table VI.
Adaptation Scenarios. We conduct experiments under three adaptation scenarios: 1) Episodic, the model parameters will be reset immediately after each optimization step of a test sample or batch; 2) Single-domain, the model is online adapted through the entire evaluation of one given test dataset (e.g., gaussian noise level 5 of ImageNet-C). Once the adaptation on this dataset is finished, the model parameters will be reset; 3) Lifelong, the model is online adapted and the parameters will never be reset (as shown in Figure 3 (Right)), which is more challenging but practical.
Evaluation Metrics. 1) Clean accuracy/error (%) on original in-distribution (ID) test samples, e.g., the original test images of ImageNet. Note that we measure the clean accuracy of all methods via (re)adaptation; 2) Out-of-distribution (OOD) accuracy/error (%) on OOD test samples, e.g., the corruption accuracy on ImageNet-C; 3) Expected Calibration Error (ECE) [70] on OOD test samples, which measures the average discrepancies between model’s confidence and accuracy within multiple confidence intervals; 4) The number of forward and backward passes during the entire TTA process. Note that the fewer #forwards and #backwards indicate less computation, leading to higher efficiency.
TABLE IV: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC lifelong test-time adaptation scenario. The model is continually adapted to the four adverse conditions for ten rounds without model reset. All results are evaluated based on the Segformer-B5 architecture. Following [64], we only show the results of the first, fourth, seventh, and last rounds due to page limits. Full results can be found in the supplementary material.
Round 1 4 7 10 All
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean
Source 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 56.7
BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0${}_{(-4.7)}$
Tent (lifelong) 69.0 40.2 60.0 57.3 66.6 36.6 58.9 54.2 64.6 33.4 55.9 51.6 62.5 30.4 52.6 48.7 52.7${}_{(-4.0)}$
CoTTA 70.9 41.1 62.4 59.7 70.8 40.6 62.6 59.7 70.8 40.5 62.6 59.7 70.8 40.5 62.6 59.7 58.4${}_{(+1.7)}$
EATA 69.1 40.5 59.8 58.1 69.3 41.8 60.1 58.6 68.8 42.5 59.4 57.9 67.9 42.8 57.7 56.3 57.0${}_{(+0.3)}$
EATA-C 71.0 44.3 63.1 61.1 72.0 47.3 64.9 63.8 71.8 48.2 64.2 64.2 72.0 48.7 64.3 64.1 61.6${}_{(+4.9)}$
Implementation Details. For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64. In EATA and ETA, the learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA). In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively. The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC. For both EATA and EATA-C, we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn. (11). More details are put in Supplementary.
5.1 Comparisons w.r.t. OOD Performance, Efficiency and Calibration Error
Results on ImageNet-C. From the results, our EATA and EATA-C significantly outperform the considered methods over both ResNet-50 and ViTBase regarding the classification accuracy, suggesting our effectiveness. With our sample-adaptive entropy loss, EATA achieves a large performance gain over the counterpart Tent (e.g., $33.4\%\rightarrow 50.5\%$ on Gaussian noise with ViTBase), verifying that removing samples with unreliable gradients and tackling samples differently benefits the test-time adaptation. More critically, EATA-C outperforms CoTTA and SAR consistently in all 15 corruption types (while EATA fails to achieve this), demonstrating the effectiveness to reduce model uncertainty with our consistency loss. Note that besides achieving strong OOD performance, EATA also alleviates the forgetting on ID samples (see Figure 3), showing the effectiveness of our anti-forgetting regularization without limiting the learning ability during adaptation (see also Table VI for ablation).
As for efficiency, the required average backward number of our EATA is 29,721 on ResNet-50, which is much fewer than the method that needs multiple data augmentations (i.e., MEMO is 50,000$\times$64) or that requires multiple rounds of optimizations (e.g., SAR is 68,145 on ResNet-50). Compared with Tent, EATA reduces the average #backward (e.g., 50,000$\rightarrow$29,721 on ResNet-50) by excluding samples with high prediction entropy and samples that are similar out of test-time optimization. In this sense, our method only needs to adapt for partial samples, resulting in higher efficiency. While EATA-C requires more forward passes, its forward passes on all test samples using the full network are gradient-free, and the more lightweight sub-network’s forward/backward passes are conducted only on the selected reliable and non-redundant samples, maintaining comparable efficiency. Note that although optimization-free methods (such as BN adaptation) do not need backward updates, their OOD performances are limited.
In terms of calibration, existing methods consistently exhibit severe overconfidence (e.g., MEMO and CoTTA are 35.63% and 27.59% on ECE with ViTBase), suggesting that overconfidence is a prevalent issue in the unsupervised test-time adaptation. Note that by filtering unreliable samples to mitigate error accumulation, EATA achieves better calibration compared with Tent (e.g., $11.7\%\rightarrow 10.5\%$ on Gaussian noise with ResNet-50). However, the miscalibration is yet significant. By further differentiating the reducible model uncertainty and the inherent data uncertainty in adaptation, our EATA-C is able to reduce the calibration error by relatively 59.8% on ResNet-50 and 64.9% on ViTBase compared with EATA, demonstrating the promising calibration effect of EATA-C on different datasets and various architectures. It is worth mentioning that while EATA-C enhances performance and efficiency, EATA-C further achieves consistent and substantial improvement over accuracy, efficiency, and calibration error compared with state-of-the-art methods, suggesting our superiority.
Results on ImageNet-R. From Table III, our EATA consistently achieves better performance than compared methods on both ResNet-50 and ViTBase, while significantly reducing the required backward propagation (e.g., 30,000normal-$\rightarrow$5,417 on ResNet-50 over Tent), demonstrating the effectiveness of adapting on partial samples during testing. Our EATA-C further improves performance by a large margin (+6.0% over EATA on ViTBase), while achieving the best calibration. More critically, EATA-C is the only method that reduces calibration error on both ResNet-50 and ViTBase, suggesting the superiority in managing model uncertainty and data uncertainty differently for calibrated test-time adaptation.
Results on CityScapes-to-ACDC. Following [64], we evaluate our method on semantic segmentation task in the lifelong adaptation scenario. From Table IV, EATA achieves comparative performance by filtering unreliable predictions and constraining the important weights from drastic change, while its counterpart, Tent, severely suffers from error accumulation. Moreover, by differentiating model uncertainty and data uncertainty in test-time adaptation, EATA-C achieves state-of-the-art performance, outperforming EATA by 4.6% and CoTTA by 3.2% in mIoU. More importantly, EATA-C can progressively accumulate knowledge over adaptation to improve its performance. In the first and tenth rounds of adaptation, EATA-C achieves 59.8% vs. 62.3% in average mIoU on the four adverse datasets, further verifying the effectiveness of our method.
5.2 Demonstration of Preventing Forgetting
Figure 3: Comparison of preventing forgetting on ImageNet-C (severity level 5) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the ID clean accuracy (after OOD adaptation). In Left, the model parameters of both Tent and our EATA are reset before adapting to a new corruption type. In Right, the model performs lifelong adaptation and the parameters will never be reset, namely Tent (lifelong) and our EATA (lifelong). The upper bound clean accuracy is estimated with the source model without adaptation on corrupted OOD data, which does not suffer from forgetting. EATA achieves higher OOD accuracy and meanwhile maintains the ID clean accuracy.
In this section, we investigate the ability of our EATA in preventing ID forgetting during test-time adaptation. The experiments are conducted on ImageNet-C with ResNet-50. We measure the anti-forgetting ability by comparing the model’s clean accuracy (i.e., on original validation data of ImageNet) before and after adaptation. To this end, we first perform test-time adaptation on a given OOD test set, and then evaluate the clean accuracy of the updated model. Here, we consider two adaptation scenarios: the single-domain adaptation, and the lifelong adaptation. We report the results of severity level 5 in Figure 3 and put the results of severity levels 1-4 into Supplementary.
From Figure 3, our EATA consistently outperforms Tent regarding the OOD corruption accuracy and meanwhile maintains the clean accuracy on ID data (in both two adaptation scenarios), demonstrating our effectiveness. It is worth noting that the performance degradation in lifelong adaptation scenario is much more severe (see Figure 3 Right). More critically, in lifelong adaptation, both the clean and corruption accuracy of Tent decreases rapidly (until degrades to 0%) after adaptation of the first three corruption types, showing that Tent in lifelong adaptation is not stable enough. In contrast, during the whole lifelong adaptation process, our EATA achieves good corruption accuracy and the clean accuracy is also very close to the clean accuracy of the model without any OOD adaptation (i.e., original clean accuracy, tested using Tent). These results demonstrate the superiority of our anti-forgetting Fisher regularizer in terms of overcoming the forgetting on ID data.
5.3 Ablation Studies
Effect of Components in $S({\bf x})$ (Eqn. 8). Our EATA accelerates test-time adaptation by excluding two types of samples out of optimization: 1) samples with high prediction entropy values (Eqn. 3) and 2) samples that are similar (Eqn. 8). We ablate both of them in Table V. Compared with the baseline $S({\bf x})\small{=}1$ (the same as Tent), introducing $S^{ent}({\bf x})$ in Eqn. (3) achieves better accuracy and fewer backwards (e.g., 49.6% (37,636) vs. 33.4% (50,000) on level 5). This verifies our motivation in Figure 2 that some high-entropy samples may hurt the performance since their gradients are unreliable. When further removing some redundant samples that are similar (Eqn. 8), our EATA further reduces the number of back-propagation (e.g., 37,636$\rightarrow$28,168 on level 5) and achieves comparable OOD error (e.g., 50.4% vs. 49.6%), demonstrating the effectiveness of our sample-efficient optimization strategy.
Effect of Components in EATA-C. By further differentiating between model uncertainty and data uncertainty based on EATA, our EATA-C achieves a substantial improvement in performance, efficiency and calibration. We conduct ablation to verify their effectiveness in Table VI. Note that adapting with simply the consistency loss in Eqn. (12) already leads to a strong baseline. By re-calibrating confidence on different samples via Eqn. (14), we further mitigate the calibration error, e.g., 3.3% (Exp1$\rightarrow$2.5(Exp3) on ECE. Compared with Exp3, Exp6 maintains comparable calibration while improving accuracy (i.e., 65.9$\rightarrow$66.4%), demonstrating that our anti-forgetting regularizer does not limit OOD adaptation. When further introducing the sample selection in Eqn. (17), EATA-C is able to effectively reduce both forward passes (i.e., 50,000$\times$2$\rightarrow$82,364) and backward passes (50,000$\rightarrow$32,364) by a large margin. Note that although accuracy and calibration are slightly compromised in EATA-C, due to the robustness of our consistency loss on corrupted images compared with Tent (see Figure 7), its improvement over the source model (65.9% vs.39.8% on accuracy, and 3.3% vs.7.5% on ECE) is still significant.
Figure 4: Effect of different entropy margins $E_{0}$ in Eqn. (3). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
Figure 5: Effect of #samples for calculating Fisher in Eqn. (11). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
Figure 4: Effect of different entropy margins $E_{0}$ in Eqn. (3). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
Figure 5: Effect of #samples for calculating Fisher in Eqn. (11). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
Figure 6: Effect of different stochastic depth ratios in EATA-C. Results obtained on ImageNet-C(Gaussian, level 5) with ViTBase.
TABLE V: Effectiveness of components in sample-adaptive weight $S({\bf x})$ in EATA on ImageNet-C (Gaussian noise) with ResNet-50.
Method Level 3 Level 5
Acc. (%) #Backwards Acc. (%) #Backwards
Baseline ($S({\bf x})\small{=}1$) 68.8 50,000 33.4 50,000
  $+S^{ent}{({\bf x})}$ (Eqn. 3) 70.7 45,302 49.6 37,636
  $+S{({\bf x})}$ (Eqn. 8) 70.8 36,057 50.4 28,168
Entropy Constant $E_{0}$ in Eqn. (3). We evaluate our EATA with different $E_{0}$, selected from {0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.55}×$\times\ln 10^{3}$, where $10^{3}$ is the class number of ImageNet. From Figure 6, our EATA achieves excellent performance when $E_{0}$ belongs to $[0.4,0.5]$. Either a smaller or larger $E_{0}$ would hamper the performance. The reasons are mainly as follows. When $E_{0}$ is small, EATA removes too many samples during adaptation and thus is unable to learn enough knowledge from the remaining samples. When $E_{0}$ is too large, some high-entropy samples would take part but contribute unreliable and harmful gradients, resulting in performance degradation. As larger $E_{0}$ leads to more backward passes, we set $E_{0}$ to 0.4×$\times\ln 10^{3}$ for efficiency-performance trade-off and fix the proportion of 0.4 for all other ImageNet experiments.
TABLE VI: Effects of components in EATA-C. Results obtained on 15 corruptions of ImageNet-C (severity level 5) with VitBase in single-domain adaptation scenario, i.e., the model parameters are reset before adapting to a new corruption. CL denotes consistency loss. ER denotes min-max entropy regularization. FR denotes anti-forgetting Fisher regularization. SS denotes active sample selection.
Experiment CL ER FR SS Average
Acc. ECE #Forwards #Backwards
1 65.6 3.3 50,000$\times$2 50,000
2 53.4 4.1 50,000$\times$2 50,000
3 65.9 2.5 50,000$\times$2 50,000
4 66.0 3.3 50,000$\times$2 50,000
5 53.0 4.5 50,000$\times$2 50,000
6 66.4 2.6 50,000$\times$2 50,000
7 65.6 4.4 83,182 33,182
8 52.9 3.4 73,323 23,323
9 (ETA-C) 65.3 3.4 82,015 32,015
10 66.2 4.3 83,429 33,429
11 53.2 3.4 73,723 23,723
12 (EATA-C) 65.9 3.3 82,364 32,364
Number of Samples for Calculating Fisher in Eqn. (11). As described in Section 4.2, the calculation of Fisher information involves a small set of unlabeled ID samples, which can be collected via existing OOD detection techniques [57]. Here, we investigate the effect of #samples $Q$, selected from {200, 300, 500, 700, 1000, 1500, 2000, 3000}. From Figure 6, our EATA achieves stable performance with $Q\geq 300$, i.e., compared with ETA, the OOD performance is comparable and the clean accuracy is much higher. These results show that our EATA does not need to collect too many ID samples, which are easy to obtain in practice.
Stochastic Depth Ratio for Obtaining Sub-Network. In Eqn. (12), We generate an extra prediction from the sub-network to measure model uncertainty, where the sub-network is obtained via stochastic depth [63] throughout the experiments. We evaluate the effect of stochastic depth ratio selected from {0.05, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35 ,0.4}. As shown in Figure 6, our EATA-C achieves a satisfying performance-calibration trade-off when the ratio belongs to $[0.15,0.25]$, where the full network consistently outperforms the sub-network while the sub-network retains sufficient capacity for learning. We fix the ratio to 0.2 for all other ImageNet experiments.
5.4 More Discussions
Results on Single Sample Adaptation (Batch Size $B=1$). In our proposed methods, we conduct test-time learning over a batch of test samples each time. When batch size $B$ equals 1, directly applying Tent and our ETA on models with batch normalization layers may fail (see Table VII), but it is not a very significant issue in practice. Actually, one can address this by either maintaining a sliding window $\{{\bf x}_{i}\}_{i=t-L}^{t}$ including $L$ previous samples and the current test sample ${\bf x}_{t}$, or data augmentation techniques adopted in existing single sample TTA methods ([14, 33]). The first way is simple to implement and incurs only minor extra costs. From Table VII, our ETA with sliding window (ETA-wnd) works well and consistently outperforms the Tent counterpart.
Wall-clock Time Speed-up of EATA and EATA-C. In the current PyTorch version, gradient computation is conducted on the full mini-batch, even if instance-wise masks are applied. To benefit from active sample selection, in EATA, we perform one forward-only pass with $B$ samples, and one extra forward and backward pass with $B^{\prime}$ selected samples. We achieve wall-time speed-up when setting $E_{0}$ in Eqn. (3) to $0.1\times\ln C$, where $C$ is the number of task classes. For ViTBase on ImageNet-C (Gaussian noise, level=5) with one Titan Xp GPU, the actual run time is 518s for Tent (33.3% accuracy), 346s for EATA (39.5% accuracy), and 402s for EATA-C (51.3% accuracy). It’s also worth mentioning that in EATA-C, the sub-network’s computation is not accelerated even if we select partial samples for each layer due to a similar setback. Actually, this is a more engineering problem, and an ideal implementation should further speed up both EATA and EATA-C.
Additional Memory by Fisher Regularizer. Since we only regularize the affine parameters of normalization layers, EATA needs very little extra memory. For ResNet-50 on ImageNet-C, the extra GPU memory at run time is only 9.8 MB, which is much less than that of Tent with batch size 64 (5,675 MB).
Performance under Mixed-and-Shifted Distributions. We evaluate Tent and our EATA/EATA-C on mixed ImageNet-C (level=3 or 5) that consists of 15 different corruption types/distribution shifts (totaling 750k images). Results in Table VIII show the stability of EATA and EATA-C in large-scale and complex TTA scenarios.
TABLE VII: Effectiveness of ETA under sliding window strategy (with different window length $L$) for single sample adaptation with ResNet-50. We report corruption accuracy (%) on ImageNet-C (Gaussian noise, level 5).
Base ResNet-50 Tent ETA Tent-wnd ($L\small{=}32)$ ETA-wnd ($L\small{=}32$) Tent-wnd ($L\small{=}64$) ETA-wnd ($L\small{=}64$)
2.2 0.1 0.1 28.1 30.8 29.5 32.4
TABLE VIII: Comparison with Tent [12] w.r.t. corruption accuracy (%) with mixture of 15 corruption types on ImageNet-C with ViTBase.
 Severity Source Tent EATA (ours) EATA-C (ours)
Level=3 63.0 69.8${}_{(+6.8)}$ 71.1${}_{(+8.1)}$ 72.4${}_{(+9.4)}$
Level=5 39.8 47.0${}_{(+7.2)}$ 58.2${}_{(+18.4)}$ 60.4${}_{(+20.6)}$
(a) Performance and calibration throughout adaptation
(b) Consistency maximization via Eqn. (12)
(c) Entropy minimization (Tent [12])
Figure 7: Comparisons of performance, calibration and stability using consistency loss in Eqn. (12) or entropy minimization loss (i.e., Tent) on ImageNet-C (Gaussian noise, severity level=5) with ViTBase. In Left, we split the dataset into 80% and 20% slices to conduct adaptation and to evaluate the adapted model. We report the performance and the calibration of the adapted model every 75 batches. In the Middle and Right, we evaluate the stability of using consistency loss and Tent under various combinations of learning rates and adaptation steps per batch. Consistency loss achieves substantially higher OOD performance and better stability while maintaining lower ECE.
Advantage of Consistency Loss over Tent [12]. We conduct a comprehensive comparison of the performance, calibration and stability between the use of consistency loss, as defined in Eqn. (12), and entropy minimization loss (i.e., Tent [12]) to reduce uncertainty at test time. The experiments are conducted on ImageNet-C (Gaussian noise, level 5) with ViTBase.
In Figure 7(a), we partition the dataset into 80% and 20% for adaptation and evaluation, respectively, to show how adaptation impacts performance and calibration over time. Analyzing Figure 7(a), we discern several key findings: 1) consistency loss consistently demonstrates superior performance and calibration throughout adaptation. 2) consistency loss is more sample-efficient, where adapting with as few as 75 batches can significantly outperform Tent [12] that adapts with 300 batches. 3) Tent [12] shows rapid degradation in performance and calibration after convergence, which indicates that it tends to over-fit. In contrast, consistency loss maintains stable performance and calibration after convergence and continuously exhibits strong generalization throughout adaptation.
In Figures 7(b) and 7(c), we study the stability of consistency loss and Tent [12] across various combinations of learning rates and adaptation steps following [71]. As illustrated in Figure 7(b), consistency loss demonstrates remarkable performance stability and benefits from the increment of adaptation steps (e.g., $55.8\%\rightarrow 58.6\%$, the best performance under 1 and 5 adaptation steps, respectively). In contrast, Tent [12] is sensitive to the combination of learning rate and adaptation steps, where its performance may deteriorate to as low as 1%. Such a phenomenon further highlights Tent [12]’s tendency to over-fit. In this sense, consistency loss emerges as the superior choice considering performance, calibration, stability and sample efficiency.
6 Conclusion
In this paper, we have proposed an Efficient Anti-forgetting Test-time Adaptation method (EATA), to improve the performance of pre-trained models on a potentially shifted test domain. To be specific, we devise a sample-efficient entropy minimization strategy that selectively performs test-time optimization with reliable and non-redundant samples. This improves the adaptation efficiency and meanwhile boosts the out-of-distribution performance. In addition, we introduce a Fisher-based anti-forgetting regularizer into test-time adaptation. With this loss, a model can be adapted continually without performance degradation on in-distribution test samples. Moreover, we design EATA with Calibration (EATA-C) for test-time adapted model’s confidence calibration. To this end, we present a consistency loss for calibrated model uncertainty reduction and a sample-aware min-max entropy regularization for confidence re-calibration, which improves the performance and calibration of test-time adaptation. Extensive experimental results on image classification and semantic segmentation demonstrate the effectiveness of our proposed methods.
Supplementary Materials for “Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting”
In the supplementary, we provide more implementation details and more experimental results of our EATA. We organize our supplementary as follows.
  • In Section A, we provide more details of our proposed EATA and EATA-C.
  • In Section B, we show more experimental results to compare the out-of-distribution performance, calibration and efficiency with state-of-the-art methods on ImageNet-C with different corruption types.
  • In Section C, we give more experimental results to demonstrate the anti-forgetting ability of our EATA.
  • In Section D, we provide more discussions on related training-time robustification studies.
Appendix A More Details of EATA and EATA-C
A.1 Overall Design of EATA
We provide an overview of our EATA. As shown in Figure A, given a trained base model $f_{\Theta^{o}}$, we perform test-time adaptation with a model $f_{\Theta}$ that is initialized from $\Theta^{o}$. During the adaptation process, we only update the parameters of batch normalization layers in $f_{\Theta}$ and froze the rest parameters. When a batch of test sample ${\mathcal{X}}\small{=}\{{\bf x}_{b}\}_{b=1}^{B}$ comes, we calculate a sample-adaptive weight $S({\bf x})$ for each test sample to identify whether the sample is active for adaptation or not. We only perform backward propagation with the samples whose $S({\bf x})\neq 0$. Moreover, we propose an anti-forgetting regularizer to prevent the model parameters $\Theta$ changing too much from $\Theta^{o}$.
Figure A: An illustration of the proposed EATA, which consists of a sample-efficient entropy minimization loss for test-time adaptation, and an anti-forgetting regularization to constrain important parameters from drastic change.
A.2 More Details on Datasets
Following the settings of Tent [12] and MEMO [14], we conduct experiments on four benchmark datasets for out-of-distribution generalization, i.e., CIFAR-10-C, ImageNet-C [9], ImageNet-R [65], and ACDC [66].
CIFAR-10-C and ImageNet-C consist of corrupted versions of the validation images on CIFAR-10 [72] and ImageNet [73], respectively. The corruptions include 15 diverse types of 4 main categories (i.e., noise, blur, weather, and digital). Each corruption type has 5 different levels of severity.
ImageNet-R contains 30,000 images with various artistic renditions of 200 ImageNet classes, which are primarily collected from Flickr and filtered by Amazon MTurk annotators.
ACDC contains four categories of images collected in adverse conditions, including fog, night, rain, and snow. Following [64], we use 400 unlabeled images from each adverse condition for adaptation.
A.3 More Experimental Protocols
Our EATA and EATA-C.
Following Tent [12] and CoTTA [64], we use ResNet-26 for CIFAR-10 experiments, ResNet-50 and ViTBase for ImageNet experiments, and Segformer-B5 for ACDC experiments. In classification experiments, the models are trained on the original CIFAR-10 or ImageNet training set and then tested on clean or the aforementioned OOD test sets. In semantic segmentation experiments, the model is trained on Cityscapes and then tested on ACDC. For a fair comparison, the parameters of ViTBase and Segformer-B5 are directly obtained from the timm [^1] https://github.com/pprp/timm and Segformer [68], respectively. ResNet-50 is trained via the official code in the torchvision [^2] https://github.com/pytorch/vision library with stochastic depth. ResNet-26 is trained via the official code of TTT by the same hyper-parameters, replacing the group norm with the batch norm, and removing the rotation head. For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64. In EATA and ETA, the learning rate is set to 0.005 on CIFAR-10, 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA). The entropy constant $E_{0}$ in Eqn. (3) is set to $0.4\times\ln C$ in classification experiments and $0.1\times\ln C$ in segmentation, where $C$ is the number of task classes. The $\epsilon$ in Eqn. (8) is set to 0.4 for CIFAR-10, and 0.05 for ImageNet and ACDC. The trade-off parameter $\beta$ in Eqn. (10) is set to 1/2,000/500 for CIFAR-10/ImageNet/ACDC to make two losses have a similar magnitude. In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively. The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC. The entropy constant $E_{0}$ in Eqn. (3) is set to 0.4/0.5×$\times\ln C$, and the $\epsilon$ in Eqn. (8) is set to 0.05/0.07 for ViTBase/ResNet-50 on ImageNet. In Eqn. (16), the trade-off parameter $\alpha$ is set to 50/500 on ImageNet/ACDC and $\beta$ is set to 0.1 on ImageNet. Note that for the semantic segmentation task, our EATA-C does not perform min-max entropy regularization and active sample selection due to the long-tailed problem [74, 75]. For both EATA and EATA-C, the moving average factor $\alpha$ in Eqn. (6) is set to 0.1, and we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn. (11).
Compared Methods. For TTA [76], BN adaptation [32], MEMO [14], CoTTA [64], and SAR [18] the hyper-parameters follow their original papers or MEMO. Specifically, the augmentation size of TTA [76] is set to 32 and 64 for CIFAR-10 and ImageNet, respectively. For BN adaptation [32], both the batch size $B$ and prior strength $N$ are set to 64. The learning rate in CoTTA [64] for ViTBase is set to 0.005, and the augmentation threshold $p_{th}$ is set to 0.1. Other hyper-parameter settings of CoTTA, and the hyper-parameter settings of MEMO [14] and SAR [18] can be found in their original paper. For Tent [12], we use SGD as the update rule with a momentum of 0.9. The batch size is 64 for both ImageNet and CIFAR-10 experiments. The learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet and 0.005 on CIFAR-10, respectively. Note that the hyper-parameters of Tent are totally the same as our EATA for a fair comparison. For TTT [11], we strictly follow their original settings except for the augmentation size at test time for ImageNet experiments. According to TTT’s implementation, the augmentation size is set to 64, which, however, is very time-consuming (e.g., about 12 GPU hours on a single Tesla V100 GPU on ImageNet-C with a specific corruption type and severity level). In our implementation, we decrease this augmentation size to 20, which has only a slight performance difference compared with augmentation size 64. For example, the performances of TTT on ImageNet-C (Gaussian noise, severity level 5) with ResNet-18 are 26.2% (64) vs. 26.0% (20). We recommend the original papers of the above methods to readers for more implementation details.
Appendix B More Results on Out-of-distribution Performance, Efficiency and Calibration
In Table D, we provide more results to compare our EATA and EATA-C with state-of-the-art methods in the single-domain adaptation scenario on ImageNet-C with the severity level 5. Although the compared methods exhibit stronger performance in the less challenging adaptation scenario, our EATA and EATA-C still demonstrate substantial superiority. Specifically, by filtering the unreliable and redundant samples from adaptation, our EATA achieves significant accuracy improvement while requiring less backward-propagation (e.g., 55.5% (72,446) vs.61.8% (32,524) over the best method SAR on ViTBase). Meanwhile, EATA-C further enhances performance while mitigating calibration error on both ResNet-50 and ViTBase, verifying the promising effectiveness of our EATA and EATA-C across different adaptation scenarios.
In Table E, we provide the complete results for the lifelong test-time adaptation on ACDC. Compared with Tent, our EATA effectively alleviates error accumulation by excluding unreliable gradients and constraining important weights from drastic changes. Furthermore, besides maintaining strong performance in the long term, our EATA-C is able to further accumulate knowledge to improve performance over adaptation, demonstrating strong potential.
In Figure B, we show the number of backward propagation of our ETA on ImageNet-C with different corruption types and severity levels. Across various corruption types, our ETA shows great superiority over existing methods in terms of adaptation efficiency. Compared with MEMO (50,000$\times$64) and Tent (50,000), our ETA only requires 31,741 backward passes (averaged over 15 corruption types) when the severity level is set to 3. The reason is that we exclude some unreliable and redundant test samples out of test-time optimization. In this case, we only need to perform backward computation on those remaining test samples, leading to improved efficiency.
In Table F, we compared ETA and EATA with existing methods on CIFAR-10-C. Under the same base model (ResNet-26 with batch normalization), ETA achieves a lower average error than Tent (19.4% vs. 20.2%) with fewer requirements of back-propagation (8,192 vs. 10,000). Moreover, the performance gain over the base model of ETA is larger than that of TTT, i.e., +9.0% vs. +7.2% on accuracy. These results are consistent with the ones on ImageNet-C/R, further demonstrating the effectiveness and superiority of our method.
Comparison with Tent using Different Learning Rates. In our sample-adaptive weight $S{({\bf x})}$ in Eqn. (8), each test sample has a specific weight $S{({\bf x})}$ and the value of $S{({\bf x})}$ is always larger than 1. In this sense, training with sample-adaptive weight $S{({\bf x})}$ indeed has the same effect as training with larger learning rates. Therefore, we compare our EATA with the baseline (Tent) using different learning rates. We increase the learning rate from $2.5\times 10^{-4}$ (which is the default of Tent) to $25.0\times 10^{-4}$ and report results in Table A.
With the learning rate increasing from $2.5\times 10^{-4}$ to $10.0\times 10^{-4}$, the error of Tent decreases from 45.3% to 43.9%, indicating that a larger learning rate may enhance the performance in some cases. However, when the learning rate becomes larger to $20.0\times 10^{-4}$, the performance of Tent degrades. More critically, our EATA method outperforms Tent with varying learning rates. These results verify that simply enlarging the learning rate is not able to achieve competitive performance with our proposed sample-adaptive adaptation method, demonstrating our superiority.
TABLE A: Comparison with Tent under different learning rates (×$\times 10^{-4}$) on ImageNet-C (Gaussian noise) regarding Error (%).
EATA (ours) Tent [12]
Severity $lr=2.5$ $lr=2.5$ $lr=10.0$ $lr=20.0$ $lr=25.0$
Level 3 42.6 45.3 43.9 44.4 45.1
Level 5 65.0 71.6 72.2 83.6 87.1
EATA under Different Random Orders. Table B records EATA’s performance (mean & stdev.) on randomly shuffled test samples with 10 different random seeds (from 2020 to 2029). From the results, EATA performs consistently across different random orders, showing the stability of EATA.
TABLE B: The mean and stdev. of corruption accuracy (%) of EATA over 10 random orders, on ImageNet-C (level 5) with ResNet-50.
Gauss. Shot Impul. Defoc. Glass Motion Zoom Avg.
34.9±${}_{\pm 0.2}$ 36.9±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$ 33.6±${}_{\pm 0.3}$ 33.3±${}_{\pm 0.2}$ 47.2±${}_{\pm 0.3}$ 52.7±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$
EATA with Different Numbers of Test Samples. We investigate the effect of the total number of test samples ($N$) in EATA, where fewer samples are sampled from the entire test set. From Table C, EATA works well and consistently outperforms Tent counterpart, regardless of the number of test samples. Meanwhile, if there are many test samples, EATA would benefit more, i.e., larger accuracy gain.
TABLE C: Comparison with Tent [12] w.r.t. corruption accuracy (%) with fewer ($N$) test samples on ImageNet-C (Gaussian noise, level 5) with ResNet-50.
Method $N\small{=}$256 $N\small{=}$512 $N\small{=}$1,024 $N\small{=}$2,048 $N\small{=}$4,096 $N\small{=}$10,000
Tent 13.7 18.2 14.8 15.2 17.5 20.9
EATA 14.8 20.1 16.3 18.7 23.5 27.6
Appendix C More Results on Prevent Forgetting
In this section, we provide more results to demonstrate the effectiveness of our EATA in preventing forgetting. We report the comparison results of EATA (lifelong) vs. Tent (lifelong) and EATA vs. Tent in Figures C and D, respectively. In the lifelong adaptation scenario, Tent suffers more severe ID performance degradation than that of reset adaptation (i.e., Figure D), showing that the more optimization steps, the more severe forgetting. Moreover, with the increase of the severity level, the ID clean accuracy degradation of Tent increases accordingly. This result indicates that the OOD adaptation with more severe distribution shifts will result in more severe forgetting. In contrast, our methods achieve higher OOD corruption accuracy and, meanwhile maintain the ID clean accuracy (competitive to the original accuracy that is tested before any OOD adaptation) in both two adaptation scenarios (reset and lifelong). These results are consistent with that in the main paper and further demonstrate the effectiveness of our proposed anti-forgetting weight regularization.
Appendix D More Discussions on Related Training-Time Robustification
To defend against distribution shifts, many prior studies seek to enlarge the training data distribution to enable it to cover the possible shift that might be encountered at test time, such as adversarial training strategies [77, 78, 79], various data augmentation techniques [80, 81, 82, 65, 83] and searching/enhancing sub-networks of a deep model [84, 85]. However, it is hard to anticipate all possible test shifts at training time. In contrast, we seek to conquer this test distribution shift by directly learning from test data.
TABLE D: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$). “BN” and “LN” denote batch and layer normalization. All results are evaluated in the single-domain adaptation scenario (i.e., the model parameters are reset before adapting to a new corruption type) except for MEMO [14]. We use * to denote episodic adaptation.
Noise Blur Weather Digital Average
Model Method Metric Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel JPEG Avg. # Forwards # Backwards
R-50 (BN) Source Acc. 1.8 3.0 1.7 18.2 10.1 13.4 20.8 14.0 22.1 21.9 58.7 5.3 17.6 22.1 37.5 17.88 50,000 0
ECE 16.6 16.1 15.9 1.8 10.7 10.7 14.7 25.3 12.9 16.7 2.3 6.7 22.9 10.5 6.0 12.64
BN Adapt Acc. 15.8 16.7 15.3 18.7 19.3 29.8 41.7 35.8 35.0 50.5 65.9 18.1 49.3 51.7 42.0 33.70 50,000 0
ECE 1.1 0.8 1.0 3.0 1.3 0.8 3.4 1.1 1.0 5.4 1.4 7.6 4.3 3.8 4.8 2.72
Tent Acc. 28.0 30.1 28.1 29.9 29.5 42.2 49.7 46.2 41.5 57.7 67.1 30.0 55.7 58.3 52.5 43.11 50,000 50,000
ECE 11.7 11.2 11.1 12.6 12.3 7.7 5.4 6.5 8.8 3.4 2.9 21.9 3.5 3.7 4.0 8.46
MEMO* Acc. 6.8 8.5 7.5 20.5 13.4 19.8 25.8 22.1 27.7 27.6 60.9 11.3 24.4 32.2 37.9 23.09 50,000$\times$65 50,000$\times$64
ECE 24.1 24.2 22.9 5.3 19.3 14.8 23.4 30.4 18.7 24.6 7.2 14.9 29.4 19.3 13.6 19.47
CoTTA Acc. 19.9 20.8 19.4 17.4 22.1 36.3 47.9 41.7 39.9 55.8 66.9 22.1 54.6 57.5 47.2 37.97 151,753 50,000
ECE 3.6 4.5 3.5 10.0 7.7 5.2 3.2 3.8 4.9 1.8 3.8 16.7 2.4 2.4 1.4 5.00
SAR Acc. 29.6 31.9 28.3 30.5 30.5 43.2 50.3 47.3 43.2 58.3 67.2 38.4 55.7 58.6 53.2 44.42 78,661 52,660
ECE 3.7 4.0 3.3 7.4 6.9 3.2 1.7 1.8 2.9 1.3 1.6 8.7 1.3 1.3 1.3 3.37
EATA (Ours) Acc. 35.6 37.8 36.1 35.3 35.4 47.6 53.1 51.4 45.9 60.1 67.6 45.8 58.2 60.3 55.4 48.36
ECE 10.5 10.2 10.3 12.9 12.3 8.3 6.6 6.7 8.2 4.7 4.5 9.9 5.1 5.3 5.6 8.08 50,000 26,188
EATA-C (Ours) Acc. 37.2 39.2 38.0 35.5 35.9 48.2 52.3 51.9 46.5 60.4 67.3 48.7 58.0 60.5 56.0 49.04
ECE 7.1 6.8 6.8 9.3 8.9 4.3 3.6 3.0 4.5 1.8 1.9 4.5 2.3 2.4 2.4 4.63 82,492 32,492
ViT (LN) Source Acc. 12.9 17.6 11.7 34.4 27.7 43.7 36.2 43.4 45.4 52.8 73.3 45.5 37.9 54.7 60.2 39.84 50,000 0
ECE 14.2 11.3 15.2 2.2 8.2 6.0 10.7 7.5 8.3 5.0 2.1 2.3 12.2 4.1 2.9 7.48
Tent Acc. 33.4 42.1 41.4 48.8 45.2 54.9 48.2 55.6 55.1 64.4 75.2 62.5 51.6 65.5 65.0 53.92 50,000 50,000
ECE 24.1 10.1 11.7 8.6 10.4 7.5 12.4 8.0 8.0 4.8 2.4 5.2 10.0 4.3 4.1 8.78
MEMO* Acc. 32.2 35.1 32.6 37.5 28.5 43.3 40.1 45.2 47.0 53.9 73.3 53.3 39.6 59.5 62.8 45.60 50,000$\times$65 50,000$\times$64
ECE 33.0 32.5 33.2 36.7 45.7 41.0 47.2 39.8 38.8 33.7 20.0 26.7 48.5 30.4 27.2 35.63
CoTTA Acc. 45.6 46.8 28.1 45.3 42.0 57.6 49.8 59.5 57.6 65.2 74.9 58.2 55.5 66.3 65.7 54.54 50,000$\times$3 50,000
ECE 9.0 8.9 9.9 7.9 12.9 8.1 12.1 7.1 9.0 6.4 4.3 6.6 10.7 6.7 6.6 8.42
SAR Acc. 43.1 44.0 44.3 50.4 48.2 55.1 49.0 56.9 56.0 64.7 75.3 62.4 53.5 65.3 64.9 55.53 86,684 72,446
ECE 8.0 7.0 7.6 6.1 6.7 5.9 9.2 5.5 5.9 3.6 2.1 4.0 7.2 3.6 3.4 5.71
EATA (Ours) Acc. 50.5 51.3 51.3 55.9 56.0 60.7 58.0 64.2 62.2 70.1 77.4 66.7 63.9 70.5 68.9 61.84
ECE 10.6 10.1 10.3 9.4 8.8 7.8 9.1 6.4 7.3 5.4 2.8 6.1 6.6 4.5 4.8 7.34 50,000 32,524
EATA-C (Ours) Acc. 56.6 57.6 57.2 57.8 59.8 64.6 64.4 69.1 67.2 73.3 78.6 67.0 70.2 73.9 71.2 65.90
ECE 5.3 4.8 5.2 5.3 4.1 3.5 4.0 2.8 2.7 2.1 1.4 2.9 2.2 1.7 2.2 3.34 82,364 32,364
TABLE E: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC in lifelong test-time adaptation scenario. The model is continually adapted to the four adverse conditions for ten rounds without resetting model parameters, i.e., lifelong adaptation. All results are evaluated based on the Segformer-B5 model.
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow cont.
Round 1 2 3 4 5 cont.
Source 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 cont.
BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 cont.
TENT-continual 69.0 40.2 60.0 57.3 68.4 39.1 60.0 56.4 67.6 37.9 59.7 55.3 66.6 36.6 58.9 54.2 65.9 35.3 57.9 53.3 cont.
CoTTA 70.9 41.1 62.4 59.7 70.9 41.0 62.5 59.7 70.9 40.8 62.6 59.7 70.8 40.6 62.6 59.7 70.8 40.6 62.6 59.7 cont.
EATA 69.1 40.5 59.8 58.1 69.3 41.1 60.0 58.4 69.3 41.5 60.1 58.6 69.3 41.8 60.1 58.6 69.2 42.1 59.9 58.5 cont.
EATA-C 71.0 44.3 63.1 61.1 71.9 44.7 63.9 62.5 72.0 46.1 64.2 63.1 72.0 47.3 64.9 63.8 71.8 46.2 64.3 64.0 cont.
Round 6 7 8 9 10 Mean
Source 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 56.7
BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0
TENT-continual 65.2 34.3 56.9 52.4 64.6 33.4 55.9 51.6 63.9 32.4 54.7 50.6 63.2 31.5 53.7 49.6 52.7 30.4 52.6 48.7 52.4
CoTTA 65.2 34.3 56.9 52.4 64.6 33.4 55.9 51.6 63.9 32.4 54.7 50.6 63.2 31.5 53.7 49.6 52.7 30.4 52.6 48.7 58.4
EATA 69.0 42.3 59.7 58.2 68.8 42.5 59.4 57.9 68.6 42.7 58.9 57.4 68.3 42.8 58.4 56.9 67.9 42.8 57.7 56.3 57.0
EATA-C 72.7 47.2 64.5 63.9 71.8 48.2 64.2 64.2 71.5 48.1 64.6 64.6 71.8 48.4 64.5 64.4 72.0 48.7 64.3 64.1 61.6
TABLE F: Comparison on CIFAR-10-C. Results are evaluated in the single-domain adaptation scenario, and averaged over 15 different corruption types and 5 severity levels (totally 75). We use * to denote episodic adaptation.
Model Acc. (%) #Forwards #Backwards
ResNet-26 (GroupNorm) 77.5 10,000 0
[…]
ResNet-26 (GroupNorm)+JT 77.2 10,000 0
[…]
ResNet-26 (BatchNorm) 71.6 10,000 0
[…]
Figure B: Comparison between ETA and Tent in terms of the number of backward propagation on ImageNet-C with different corruption types and severity levels.
Figure C: Comparison of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model performs lifelong adaptation, in which the model parameters will never be reset.
Figure D: Comparisons of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50. We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation). The model parameters of both Tent and our EATA are reset before adapting to a new corruption type.

### v2

@@
 Index Terms:
-Out-of-Distribution Generalization, Test-Time Adaptation, Calibration, Catastrophic, Forgetting, Robustness.
+Out-of-Distribution Generalization, Test-Time Adaptation, Confidence Calibration, Catastrophic Forgetting.
 1 Introduction
@@
 Over-Confident Predictions.
-Existing methods like Tent [12] and SAR [18] primarily focus on reducing the model’s prediction uncertainty through unsupervised prediction entropy minimization, which enhances the model’s predicting confidence, thereby learning a decision boundary positioned in low-density sample regions [19, 20].
-However, a critical oversight of these methods is their neglect of the irreducible data uncertainty inherent in test samples.
-They encourage highly certain predictions for all test samples even in cases of underlying data uncertainty or even if the prediction is wrong, leading to overly certain (or overconfident) predictions.
-This phenomenon is particularly concerning in high-risk applications [21, 22], posing potential safety risks.
+Existing methods like Tent [12] and SAR [18] primarily rely on test-time entropy minimization for model adaptation, which greedily enhances the model’s confidence and minimizes the predictive uncertainty for test samples, without distinguishing between model-induced and data-induced uncertainties.
+Consequently, even when the input data is naturally complex or highly corrupted (i.e., with irreducible data uncertainty), the model is forced to make one-hot confident predictions where it should remain uncertain, leading to over-confident and potentially incorrect outputs.
+This phenomenon is particularly concerning in high-risk applications, such as autonomous driving [19] and medical diagnosis [20], posing potential safety risks.
 To address the efficiency and forgetting issue, we have proposed an Efficient Anti-forgetting Test-time Adaptation (EATA) method consisting of a sample-efficient optimization strategy and a weight regularizer.
@@
 With this regularization, the model can continually adapt to OOD test samples without performance degradation on ID test samples.
-To mitigate the over-confident issue, it is essential to differentiate between various origins of uncertainty in TTA, i.e., reducible model uncertainty and inherent data uncertainty.
-Inspired by the human tendency to base predictions on available knowledge and problem difficulty, rather than defaulting to high certainty in unfamiliar situations, we propose to utilize the model uncertainty for calibrated uncertainty reduction and leverage the inherent data uncertainty to directly re-calibrate predicting confidence.
-To this end, we further propose EATA with Calibration, namely EATA-C.
-To be specific, EATA-C estimates model uncertainty by measuring the divergence between predictions from the full network and its randomly sampled sub-networks.
-This allows for the reduction of model uncertainty by minimizing the divergence, promoting consistent rather than overconfident predictions during testing.
-Additionally, we develop a simple but effective indicator of data uncertainty based on the disagreement among predicted labels, which suggests whether a sample nears decision boundaries.
-Then, we devise an auxiliary min-max entropy regularizer to selectively lower confidence for inherently uncertain samples through an entropy maximization loss, while simultaneously boosting confidence for certain samples via an entropy minimization loss.
-EATA-C differs from EATA on the adaptation objective, while it still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA.
+To mitigate overconfidence, we differentiate between various origins of uncertainty in TTA: 1) Reducible model uncertainty, which arises from not knowing the optimal model parameters to describe the data due to insufficient training [22]; 2) Irreducible data uncertainty that arises from inherent noise or variability in the data, cannot be reduced by additional training [23].
+Based on their characteristics, we aim to reduce the model uncertainty at test time for domain adaptation, while accurately reflecting data uncertainty in model predictions to ensure confidence calibration.
+To this end, we further devise EATA with Calibration, namely EATA-C.
+Specifically, EATA-C estimates model uncertainty by measuring the divergence between predictions from the full network and its randomly sampled sub-networks.
+By minimizing this divergence, our EATA-C reduces model uncertainty and promotes consistent, rather than overconfident, predictions for model adaptation during testing.
+Additionally, we introduce a data uncertainty indicator based on prediction disagreement, which effectively detects ambiguous samples near decision boundaries where conflicting predictions are more likely to occur.
+We then incorporate a min-max entropy regularizer to selectively adjust the prediction confidence based on this data uncertainty estimation.
+Note that EATA-C differs from EATA on the adaptation objective, while it still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA.
 We summarize our main contributions as follows.
   • We propose an Efficient Anti-forgetting Test-time Adaptation (EATA) method.
-Specifically, we develop an active sample identification scheme to filter out non-reliable and redundant test samples from model adaptation, thereby improving adaptation efficiency by reducing backward propagation.
-Moreover, we extend the label-dependent Fisher regularizer to test samples with pseudo label generation, which prevents drastic changes in important model weights and thus alleviates the issue of model forgetting on in-distribution test samples.
-  • We further propose EATA with Calibration (EATA-C), which decouples the model and data uncertainty to achieve confidence-calibrated TTA.
-Specifically, EATA-C estimates model uncertainty using the divergence between full network and sub-network predictions, incorporating a consistency loss to reduce this uncertainty without leading to overconfidence.
-For data uncertainty, EATA-C utilizes label prediction disagreements, employing min-max entropy regularization to selectively adjust prediction confidence for effective re-calibration.
-  • We demonstrate that our proposed EATA method improves both the performance and efficiency of test-time adaptation and also alleviates the long-neglected catastrophic forgetting issue, and EATA-C further achieves better performance and calibration without additional backward computations compared with EATA.
+Specifically, we reveal that test samples contribute differently to adaptation, and develop an active sample identification scheme to filter out non-reliable and redundant test samples from adaptation, thereby improving TTA efficiency.
+Moreover, we extend the label-dependent Fisher regularizer to test samples with pseudo label generation, which prevents drastic changes in important model weights and helps alleviate the issue of model forgetting on in-distribution test samples.
+  • We further introduce EATA with Calibration (EATA-C), which differentiates between reducible and irreducible uncertainty during testing to design a calibration-driven learning objective.
+Specifically, EATA-C estimates model uncertainty using the divergence between full network and sub-network predictions, incorporating a consistency loss to reduce this uncertainty for adaptation.
+Regarding data uncertainty, EATA-C leverages prediction disagreements and applies min-max entropy regularization to selectively adjust confidence for calibration enhancement.
+  • We demonstrate that our proposed EATA method improves both the performance and efficiency of test-time adaptation and also alleviates the long-neglected catastrophic forgetting issue.
+Our EATA-C further achieves better performance and calibration, with computational and memory efficiency comparable to EATA.
+A short version of this work was published in ICML 2022 [24].
+This paper extends our preliminary version from the following aspects: 1) We explore the calibrated test-time adaptation, which aims to provide calibrated predicting confidence that reflects the true likelihood of correctness during unsupervised adaptation; 2) To solve the over-confident issue, we develop a test-time consistency loss that leverages the reducible model uncertainty for calibrated uncertainty reduction, and devise a min-max entropy regularizer to re-calibrate predicting confidence based on the inherent data uncertainty; 3) We provide analyses about the impact of different uncertainty reduction strategies, empirically verifying that our consistency loss overcomes the issue of over-fitting and over-confident in entropy minimization loss when adapting to the test data; 4) We provide extensive new empirical evaluations on image classification and semantic segmentation tasks with various model architectures, demonstrating that EATA-C achieves substantially better performance and calibration over EATA, e.g., improving accuracy by 6.5%, while reducing calibration error by relatively 64.9% on ImageNet-C dataset with ViT-Base [25].
 Figure 1: An illustration of our proposed Efficient Anti-forgetting Test-time Adaptation with Calibration (EATA-C) method.
@@
 We divide the discussion on related works based on the different adaptation settings summarized in Table I and further review existing methods for model’s uncertainty calibration.
-Test-Time Adaptation (TTA) aims to improve model accuracy on OOD test data through model adaptation with test samples, which has been successfully applied to various tasks, such as image classification [11, 12], image super-resolution [24], image captioning [25], object detection [26], video recognition [27, 28], etc.
+Test-Time Adaptation (TTA) aims to improve model accuracy on OOD test data through model adaptation with test samples.
 Existing test-time training methods, e.g., TTT [11], TTT++ [13], TTT-MAE [26], and MT3 [27], jointly train a source model via both supervised and self-supervised objectives, and then adapt the model via self-supervised objective at test time.
@@
 We empirically validate the existence of this issue in Figure 3, where the updated model has a consistently lower accuracy on ID test samples than the original model.
-Moreover, existing entropy-based test-time adaptation methods like Tent [12] and SAR [18] consistently encourage producing one-hot predictions with high confidence.
-However, in practice, incoming test samples can be severely corrupted [9], inheriting irreducible data uncertainty.
-Considering the data uncertainty, these test samples ideally should be predicted with relatively low confidence.
-Nevertheless, the data uncertainty is often overlooked by methods based on entropy minimization.
-Consequently, the adapted model often produces highly confident predictions (called overconfidence), even when the test samples are uncertain.
+Moreover, existing entropy-based test-time adaptation methods like Tent [12] and SAR [18] consistently encourage the model to produce one-hot highly confident predictions.
+However, in practice, input test samples can be naturally complex and severely corrupted [9], resulting in irreducible data uncertainty.
+Ideally, these samples should be predicted with relatively low confidence to reflect their ambiguity.
+Nevertheless, the data uncertainty is often overlooked by methods based on entropy minimization,causing the adapted model to produce highly confident predictions (called overconfidence), even when predictions should remain uncertain.
 Such misleading predictions raise potential safety concerns for real-world application scenarios.
@@
 Figure 2: Effect of different test samples in test-time entropy minimization [12].
-We adapt a model on partial samples (top $p$samples with the highest or lowest entropy values), and then evaluate the adapted model on all test samples.
+We adapt a model on partial samples (top pp% samples with the highest or lowest entropy values), and then evaluate the adapted model on all test samples.
 Results are obtained on ImageNet-C (Gaussian noise, level 3) and ResNet-50 (base accuracy is 27.6%).
@@
 4.3 Consistency-Based Uncertainty Minimization
-As mentioned in Section 4.1, EATA enhances model performance by minimizing prediction entropy.
-This strategy aims to reduce uncertainty in predictions and learn a decision boundary in the low-density regions of the test samples [58, 59].
-However, a persistent challenge with entropy minimization is its tendency to yield overly certain predictions, even in cases where the test samples themselves are inherently uncertain, such as severely corrupted images [9].
-Thus, EATA may still result in overconfident predictions that fail to accurately account for the inherent data uncertainty.
+As mentioned in Section 4.1, EATA conducts model adaptation by prediction entropy minimization.
+This strategy aims to reduce uncertainty in predictions and learn decision boundaries in low-density regions of the test samples [55, 56].
+However, a persistent limitation of entropy minimization is its tendency to yield overly certain predictions, where the model is forced to make one-hot confident predictions even for ambiguous input data with irreducible data uncertainty.
+Thus, EATA may still result in overconfident predictions that do not accurately reflect the inherent data uncertainty.
 To address this, we further propose EATA with Calibration (EATA-C), shall be depicted in Sections 4.3 and 4.4.
@@
 Here, we use only two predictions and the former is indispensable since we select it as the final prediction.
-Thus, at test time, we minimize divergence to encourage consistent predictions instead of confidence-related entropy loss, thereby alleviating the overconfidence issue.
+During TTA, we minimize this divergence to promote consistent predictions for model updates, rather than greedily increasing confidence which can result in overconfident outputs.
 Consistency Loss.
@@
 $\hat{{\bf y}}_{fuse}=($ $\hat{{\bf y}}+(1-p)\cdot\hat{{\bf y}}_{sub})/(2-p),$ (13)
-where $D_{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence [61] and $p$ is a constant for balancing.
-Here, we calculate the divergence between $\hat{{\bf y}}$ and $\hat{{\bf y}}_{fuse}$ (rather than $\hat{{\bf y}}_{sub}$), since encouraging the sub-network to achieve the same performance as the full one is relatively hard.
+where $D_{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence [58] and $p$ is a constant for smoothing.
+Here, we calculate the divergence between $\hat{{\bf y}}_{sub}$ and $\hat{{\bf y}}_{fuse}$ (rather than $\hat{{\bf y}}$), since encouraging the sub-network to achieve the same performance as the full one is relatively hard.
 Thus, inspired by label smoothing [59], we softly fuse $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ in Eqn.
@@
 To this end, we detach the gradient from $\hat{{\bf y}}$ and concentrate the optimization solely on $\hat{{\bf y}}_{sub}$.
-This strategy is designed to facilitate knowledge transfer from the full network to its sub-network, thereby enhancing the sub-network’s performance while simultaneously reducing the model uncertainty of the full network.
+This strategy is designed to facilitate knowledge transfer from the full network to its sub-network, thereby enhancing the sub-network’s performance while reducing the full networks’ model uncertainty to adapt the network to the test domain.
 Remark on Efficiency.
 Although consistency loss requires two forward passes from both the full network and the sub-network for each sample, the full network’s forward pass is gradient-free without back-propagation and the sub-network’s forward/backward pass is less computationally intensive.
-Moreover, we only perform sub-network’s forward/backward pass on the selected reliable and non-redundant samples as outlined in Algorithm 1.
-As a result, the use of consistency loss remains efficient.
+Moreover, we only perform sub-network’s forward/backward passes on the selected reliable and non-redundant samples as outlined in Algorithm 1.
+As a result, the use of consistency loss remains efficient as shown in Table VII.
 Algorithm 1 The pipeline of proposed EATA and EATA-C.
@@
 This process involves categorizing samples into two distinct groups—‘certain’ and ‘uncertain’—based on the aforementioned prediction consistency.
-This identification is achieved by comparing the predicted labels between the full network and a sub-network.
-Samples exhibiting mismatched predictions are deemed ‘uncertain’, indicating either high intrinsic data uncertainty or a high level of complexity that challenges the model’s recognition capability.
-For these uncertain samples, we aim to lower their predictive confidence.
-This is achieved by maximizing the prediction entropy, effectively acknowledging the model’s lack of confidence in these cases.
+This design is inspired by margin-based learning approaches [61, 62] which indicated that samples near decision boundaries are inherently more uncertain and have been well justified with theoretical guarantees.
+Specifically, we achieve categorization by comparing the predicted labels between the full network and a sub-network.
+Samples that exhibit mismatched predictions are deemed ‘uncertain’, suggesting their proximity to decision boundaries and high intrinsic data uncertainty.
+Note that unlike the consistency loss that measures model uncertainty, in which samples across the data space may yield low consistency loss, data uncertainty is reflected more prominently through prediction disagreements near the decision boundary (see Figure C for illustration).
+For identified uncertain samples, we aim to lower their predictive confidence by maximizing the prediction entropy, effectively acknowledging the model’s lack of confidence in these cases.
 Conversely, for samples where predictions are consistent, labeled as ‘certain’, we conduct the opposite strategy, i.e., boosting prediction confidence through entropy minimization.
@@
 $\min_{\tilde{\Theta}_{sub}}C({\bf x})E({\bf x};\Theta_{sub}),$~~~~~~~~~~~~~~~~~ (14)
-$C({\bf x})=\begin{cases}1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})=\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ -1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})\neq\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ \end{cases}$ (15)
+[…]
 where $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ denote the prediction of the full and sub-network respectively, $\Theta_{sub}$ denotes the parameters of the sub-network and $\tilde{\Theta}_{sub}\subset\Theta_{sub}$ denotes parameters involved in model adaptation.
@@
 Overall Objective.
-The methods proposed in Sections 4.3 and 4.4 are devised to address the overconfident issue in TTA, but still suffer from catastrophic forgetting when the important model weights for the in-distribution domain are modified significantly during adaptation.
-Therefore, we jointly optimize the model with the anti-forgetting regularizer and further reduce the required backward computations with the active sample selection criterion proposed in EATA.
+The methods proposed in Sections 4.3 and 4.4 are devised to address the overconfident issue in TTA, but still suffer from catastrophic forgetting when important model weights for the in-distribution domain are significantly modified during adaptation.
+Therefore, we jointly optimize the model with the anti-forgetting regularizer and further reduce the required backward computations with the active sample selection criterion in EATA.
 Then, the overall objective of EATA-C can be formulated as:
@@
 5 Experiments
-TABLE II: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$).
+TABLE II: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, $\uparrow$) and Expected Calibration Error($\downarrow$).
 “BN” and “LN” denote batch and layer normalization, respectively.
@@
 All results are evaluated under the lifelong adaptation scenario except for Tent [12] and MEMO [14], which suffer severely from error accumulation.
-We use * and ${}^{\dagger}$ to denote episodic and single-domain adaptation, respectively.
+We use * and to denote episodic and single-domain adaptation, respectively.
 Noise Blur Weather Digital Average
@@
 ECE 1.1 0.8 1.0 3.0 1.3 0.8 3.4 1.1 1.0 5.4 1.4 7.6 4.3 3.8 4.8 2.72
-${}^{\dagger}$ Acc. 28.0 30.1 28.1 29.9 29.5 42.2 49.7 46.2 41.5 57.7 67.1 30.0 55.7 58.3 52.5 43.11 50,000 50,000
+Tent Acc. 28.0 30.1 28.1 29.9 29.5 42.2 49.7 46.2 41.5 57.7 67.1 30.0 55.7 58.3 52.5 43.11 50,000 50,000
 ECE 11.7 11.2 11.1 12.6 12.3 7.7 5.4 6.5 8.8 3.4 2.9 21.9 3.5 3.7 4.0 8.46
-MEMO* Acc. 6.8 8.5 7.5 20.5 13.4 19.8 25.8 22.1 27.7 27.6 60.9 11.3 24.4 32.2 37.9 23.09 50,000$\times$65 50,000$\times$64
+[…]
 ECE 24.1 24.2 22.9 5.3 19.3 14.8 23.4 30.4 18.7 24.6 7.2 14.9 29.4 19.3 13.6 19.47
@@
 ECE 3.7 7.4 9.1 14.4 15.1 12.6 9.6 11.7 13.3 8.4 7.0 16.8 8.5 8.4 8.5 10.29
-EATA (Ours) Acc. 35.6 38.7 37.5 35.9 36.1 47.6 53.1 50.6 45.6 59.5 67.1 45.2 57.6 59.8 55.4 48.36
-ECE 10.5 13.4 14.7 18.5 18.5 14.7 12.8 14.1 16.1 11.6 10.4 18.6 13.3 13.1 13.9 14.28 50,000 29,721
-EATA-C (Ours) Acc. 37.2 40.9 39.9 36.6 37.1 48.5 52.7 51.9 47.2 60.4 67.0 49.0 57.6 60.1 56.1 49.48
-ECE 7.1 6.7 7.2 10.1 9.5 5.9 4.9 4.8 5.8 3.2 3.2 6.0 4.0 3.9 3.9 5.74 83,312 33,312
+ROID Acc. 36.7 38.6 35.9 29.1 28.8 40.6 46.5 49.5 41.8 55.6 65.3 43.5 53.6 56.1 52.8 44.95 80,380 80,380
+ECE 36.6 38.5 35.8 29.0 28.7 40.5 46.4 49.4 41.7 55.5 65.2 43.4 53.5 56.0 52.7 44.85
+RDump Acc. 35.6 34.4 36.1 32.7 34.6 45.5 51.6 50.3 44.1 59.9 66.8 46.3 56.8 59.1 54.7 47.23 50,000 26,375
+ECE 10.5 11.4 12.6 13.9 14.4 9.4 8.1 8.1 9.4 6.3 5.5 11.8 6.1 7.0 6.3 9.40
+[…]
+ECE 7.9 5.4 2.4 2.0 1.8 1.9 3.8 19.5 67.6 76.4 86.2 91.0 90.4 91.3 91.3 42.60
+Acc. 35.6 38.7 37.5 35.9 36.1 47.6 53.1 50.6 45.6 59.5 67.1 45.2 57.6 59.8 55.4 48.36
+EATA (Ours) ECE 10.5 13.4 14.7 18.5 18.5 14.7 12.8 14.1 16.1 11.6 10.4 18.6 13.3 13.1 13.9 14.28 50,000 29,721
+Acc. 37.2 40.9 39.9 36.6 37.1 48.5 52.7 51.9 47.2 60.4 67.0 49.0 57.6 60.1 56.1 49.48
+EATA-C (Ours) ECE 7.1 6.7 7.2 10.1 9.5 5.9 4.9 4.8 5.8 3.2 3.2 6.0 4.0 3.9 3.9 5.74 83,312 33,312
 ViT (LN) Source Acc. 12.9 17.6 11.7 34.4 27.7 43.7 36.2 43.4 45.4 52.8 73.3 45.5 37.9 54.7 60.2 39.84 50,000 0
 ECE 14.2 11.3 15.2 2.2 8.2 6.0 10.7 7.5 8.3 5.0 2.1 2.3 12.2 4.1 2.9 7.48
-${}^{\dagger}$ Acc. 33.4 42.1 41.4 48.8 45.2 54.9 48.2 55.6 55.1 64.4 75.2 62.5 51.6 65.5 65.0 53.92 50,000 50,000
+Tent Acc. 33.4 42.1 41.4 48.8 45.2 54.9 48.2 55.6 55.1 64.4 75.2 62.5 51.6 65.5 65.0 53.92 50,000 50,000
 ECE 24.1 10.1 11.7 8.6 10.4 7.5 12.4 8.0 8.0 4.8 2.4 5.2 10.0 4.3 4.1 8.78
-MEMO* Acc. 32.2 35.1 32.6 37.5 28.5 43.3 40.1 45.2 47.0 53.9 73.3 53.3 39.6 59.5 62.8 45.60 50,000$\times$65 50,000$\times$64
+[…]
 ECE 33.0 32.5 33.2 36.7 45.7 41.0 47.2 39.8 38.8 33.7 20.0 26.7 48.5 30.4 27.2 35.63
-CoTTA Acc. 45.6 58.8 58.4 40.1 50.9 50.4 44.4 46.1 51.4 52.2 57.7 36.5 53.9 55.8 55.4 50.51 50,000$\times$3 50,000
+[…]
 ECE 9.0 15.8 21.0 29.3 28.3 29.1 34.5 29.8 29.6 30.0 29.2 36.3 29.0 31.2 31.8 27.59
@@
 ECE 8.0 8.4 9.5 9.1 11.4 9.6 13.0 9.9 9.6 7.8 4.5 10.0 13.8 7.5 7.5 9.30
-EATA (Ours) Acc. 50.5 55.6 56.0 54.9 56.3 61.1 59.8 64.3 64.0 70.1 77.4 65.5 63.1 70.4 69.5 62.57
-ECE 10.6 13.9 15.7 16.4 17.3 15.9 17.5 14.8 15.7 12.8 9.0 15.4 17.5 13.3 13.9 14.63 50,000 36,688
-EATA-C (Ours) Acc. 56.8 60.2 59.8 58.0 60.9 65.2 65.5 69.7 67.8 74.0 78.9 66.7 70.3 74.1 71.8 66.65
-ECE 5.2 5.1 5.9 7.1 6.2 5.6 5.9 5.0 5.0 4.1 3.3 5.4 4.7 4.2 4.7 5.14 83,184 33,184
+ROID Acc. 48.8 49.5 49.0 54.1 54.2 58.8 55.7 62.7 61.3 69.7 77.0 65.5 64.1 69.6 68.3 60.56 80,739 80,739
+ECE 48.7 49.4 48.9 54.0 54.1 58.7 55.6 62.6 61.2 69.6 76.9 65.4 64.0 69.5 68.2 60.46
+RDump Acc. 50.5 48.4 51.8 54.0 55.3 59.4 55.9 63.3 60.7 70.8 76.8 66.8 60.9 69.5 68.4 60.84 50,000 32,050
+ECE 10.6 9.7 10.5 8.9 9.4 7.9 10.0 6.9 7.4 5.4 3.2 6.7 7.6 5.1 5.0 7.62
+[…]
+ECE 10.6 13.7 13.7 14.5 14.4 11.3 11.6 9.4 10.2 7.8 3.9 10.5 9.8 7.3 7.4 10.40
+Acc. 50.5 55.6 56.0 54.9 56.3 61.1 59.8 64.3 64.0 70.1 77.4 65.5 63.1 70.4 69.5 62.57
+EATA (Ours) ECE 10.6 13.9 15.7 16.4 17.3 15.9 17.5 14.8 15.7 12.8 9.0 15.4 17.5 13.3 13.9 14.63 50,000 36,688
+Acc. 56.8 60.2 59.8 58.0 60.9 65.2 65.5 69.7 67.8 74.0 78.9 66.7 70.3 74.1 71.8 66.65
+EATA-C (Ours) ECE 5.2 5.1 5.9 7.1 6.2 5.6 5.9 5.0 5.0 4.1 3.3 5.4 4.7 4.2 4.7 5.14 83,184 33,184
 TABLE III: Comparison on ImageNet-R. Results are evaluated in the single-domain adaptation scenario. We use * to denote episodic adaptation.
-Model Acc. (%) ECE (%) #Forwards #Backwards
+Model
+Acc. (%)
+ECE (%)
+#Forwards
+#Backwards
 ResNet-50(BN) 38.0 17.7 30,000 0
+$\bullet$~BN [29] 40.4(+2.4) 13.4(-4.3) 30,000 0
+$\bullet$~Tent [12] 42.3(+4.3) 17.8(+0.1) 30,000 30,000
 […]
+$\bullet$~CoTTA [63] 42.4(+4.4) 15.8(-1.9) 90,000 30,000
+$\bullet$~SAR [18] 42.7(+4.7) 14.6(-3.1) 47,755 32,877
+$\bullet$~ROID [64] 48.6(+10.5) 48.1(+30.4) 48,303 48,303
 […]
+$\bullet$~EATA (Ours) 44.9(+6.9) 16.7(-1.0) 30,000 5,417
+$\bullet$~EATA-C (Ours) 47.1(+9.1) 13.3(-4.4) 35,122 5,122
+ViT(LN) 52.5 5.0 30,000 0
+$\bullet$~Tent [12] 54.2(+1.7) 7.4(+2.4) 30,000 30,000
 […]
+$\bullet$~CoTTA [63] 56.4(+3.9) 7.4(+2.4) 90,000 30,000
+$\bullet$~SAR [18] 55.0(+2.5) 5.2(+0.2) 47,119 33,844
+$\bullet$~ROID [64] 62.2(+9.7) 61.7(+56.7) 49,795 49,795
 […]
-[…]
-[…]
-[…]
-ViT(LN) 52.5 5.0 30,000 0
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+$\bullet$~EATA (Ours) 58.2(+5.7) 5.8(+0.8) 30,000 6,053
+$\bullet$~EATA-C (Ours) 64.2(+11.7) 3.9(-1.1) 36,395 6,395
 Datasets and Models.
-We conduct experiments on three benchmarks datasets for OOD generalization, i.e., ImageNet-C [9] (contains corrupted images in 15 types of 4 main categories and each type has 5 severity levels) and ImageNet-R [65] for image classification, and ACDC [66] for semantic segmentation.
-We use ResNet-50 (R-50) [1] and VitBase (ViT) [67] for ImageNet experiments, and Segformer-B5 [68] for ACDC [66] experiments.
-The models are trained on ImageNet or CityScapes [69] training set with stochastic depth regularization [63] and then tested on clean or the above OOD test sets.
+We conduct experiments on three benchmark datasets for OOD generalization: ImageNet-C [9] (contains corrupted images in 15 types of 4 main categories and each type has 5 severity levels) and ImageNet-R [66] for image classification; and ACDC [67] for semantic segmentation.
+We use ResNet-50 (R-50) [1] and ViT-Base (ViT) [25] for ImageNet experiments, and Segformer-B5 [68] for ACDC [67] experiments.
+The models are trained on ImageNet or CityScapes [69] training set with stochastic depth regularization [60] and tested on clean or OOD test sets.
 Compared Methods.
@@
 MEMO [14] maximizes the prediction consistency of different augmented copies regarding a given test sample.
-CoTTA [64] minimizes the cross entropy between the predictions of the student network and its mean teacher at test time.
-SAR [18] selects reliable samples to perform test time sharpness-aware entropy minimization.
+SAR [18] selects reliable samples for test time sharpness-aware entropy minimization.
+CoTTA [63] and DAT [70] minimize the cross entropy between the student network and its mean teacher during testing.
+RDump [71] periodically resets model parameters based on our EATA.
+TEA [65] employs energy-based contrastive learning with negative sample generation.
+ROID [64] minimizes the diversity-weighted soft likelihood ratio loss.
 We denote EATA without weight regularization in Eqn.
@@
 Source 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 69.1 40.3 59.7 57.8 56.7
-BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0${}_{(-4.7)}$
-Tent (lifelong) 69.0 40.2 60.0 57.3 66.6 36.6 58.9 54.2 64.6 33.4 55.9 51.6 62.5 30.4 52.6 48.7 52.7${}_{(-4.0)}$
-CoTTA 70.9 41.1 62.4 59.7 70.8 40.6 62.6 59.7 70.8 40.5 62.6 59.7 70.8 40.5 62.6 59.7 58.4${}_{(+1.7)}$
-EATA 69.1 40.5 59.8 58.1 69.3 41.8 60.1 58.6 68.8 42.5 59.4 57.9 67.9 42.8 57.7 56.3 57.0${}_{(+0.3)}$
-EATA-C 71.0 44.3 63.1 61.1 72.0 47.3 64.9 63.8 71.8 48.2 64.2 64.2 72.0 48.7 64.3 64.1 61.6${}_{(+4.9)}$
-Implementation Details.
-For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64.
-In EATA and ETA, the learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA).
-In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively.
-The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC.
-For both EATA and EATA-C, we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn.
-(11).
-More details are put in Supplementary.
+BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0(-4.7)
+Tent (lifelong) 69.0 40.2 60.0 57.3 66.6 36.6 58.9 54.2 64.6 33.4 55.9 51.6 62.5 30.4 52.6 48.7 52.7(-4.0)
+CoTTA 70.9 41.1 62.4 59.7 70.8 40.6 62.6 59.7 70.8 40.5 62.6 59.7 70.8 40.5 62.6 59.7 58.4(+1.7)
+DAT 71.7 44.6 63.8 62.2 68.0 42.0 60.9 59.4 66.1 40.6 59.8 57.8 63.8 39.6 58.2 55.4 57.0(+0.3)
+EATA 69.1 40.5 59.8 58.1 69.3 41.8 60.1 58.6 68.8 42.5 59.4 57.9 67.9 42.8 57.7 56.3 57.0(+0.3)
+EATA-C 71.0 44.3 63.1 61.1 72.0 47.3 64.9 63.8 71.8 48.2 64.2 64.2 72.0 48.7 64.3 64.1 61.6(+4.9)
+[…]
 5.1 Comparisons w.r.t. OOD Performance, Efficiency and Calibration Error
 Results on ImageNet-C.
-From the results, our EATA and EATA-C significantly outperform the considered methods over both ResNet-50 and ViTBase regarding the classification accuracy, suggesting our effectiveness.
-With our sample-adaptive entropy loss, EATA achieves a large performance gain over the counterpart Tent (e.g., $33.4\%\rightarrow 50.5\%$ on Gaussian noise with ViTBase), verifying that removing samples with unreliable gradients and tackling samples differently benefits the test-time adaptation.
-More critically, EATA-C outperforms CoTTA and SAR consistently in all 15 corruption types (while EATA fails to achieve this), demonstrating the effectiveness to reduce model uncertainty with our consistency loss.
+From Table II, our EATA and EATA-C consistently surpass existing approaches regarding adaptation accuracy, e.g., the average accuracy of 48.4% (EATA) vs.
+45.0% (ROID) on ResNet-50.
+Importantly, EATA yields a remarkable performance gain over its counterpart Tent, e.g., $33.4\%\rightarrow 50.5\%$ on Gaussian Noise with ViT-Base, suggesting the significance of removing samples with unreliable gradients and tackling samples differently in the TTA process.
+Our enhanced method, EATA-C, further boosts adaptation accuracy by a large margin, which consistently outperforms TEA and ROID in all 15 corruption types over both ResNet-50 and ViT-Base, suggesting our effectiveness.
 Note that besides achieving strong OOD performance, EATA also alleviates the forgetting on ID samples (see Figure 3), showing the effectiveness of our anti-forgetting regularization without limiting the learning ability during adaptation (see also Table VI for ablation).
-As for efficiency, the required average backward number of our EATA is 29,721 on ResNet-50, which is much fewer than the method that needs multiple data augmentations (i.e., MEMO is 50,000$\times$64) or that requires multiple rounds of optimizations (e.g., SAR is 68,145 on ResNet-50).
-Compared with Tent, EATA reduces the average #backward (e.g., 50,000$\rightarrow$29,721 on ResNet-50) by excluding samples with high prediction entropy and samples that are similar out of test-time optimization.
-In this sense, our method only needs to adapt for partial samples, resulting in higher efficiency.
-While EATA-C requires more forward passes, its forward passes on all test samples using the full network are gradient-free, and the more lightweight sub-network’s forward/backward passes are conducted only on the selected reliable and non-redundant samples, maintaining comparable efficiency.
-Note that although optimization-free methods (such as BN adaptation) do not need backward updates, their OOD performances are limited.
-In terms of calibration, existing methods consistently exhibit severe overconfidence (e.g., MEMO and CoTTA are 35.63% and 27.59% on ECE with ViTBase), suggesting that overconfidence is a prevalent issue in the unsupervised test-time adaptation.
-Note that by filtering unreliable samples to mitigate error accumulation, EATA achieves better calibration compared with Tent (e.g., $11.7\%\rightarrow 10.5\%$ on Gaussian noise with ResNet-50).
-However, the miscalibration is yet significant.
-By further differentiating the reducible model uncertainty and the inherent data uncertainty in adaptation, our EATA-C is able to reduce the calibration error by relatively 59.8% on ResNet-50 and 64.9% on ViTBase compared with EATA, demonstrating the promising calibration effect of EATA-C on different datasets and various architectures.
-It is worth mentioning that while EATA-C enhances performance and efficiency, EATA-C further achieves consistent and substantial improvement over accuracy, efficiency, and calibration error compared with state-of-the-art methods, suggesting our superiority.
-Results on ImageNet-R.
-From Table III, our EATA consistently achieves better performance than compared methods on both ResNet-50 and ViTBase, while significantly reducing the required backward propagation (e.g., 30,000normal-$\rightarrow$5,417 on ResNet-50 over Tent), demonstrating the effectiveness of adapting on partial samples during testing.
-Our EATA-C further improves performance by a large margin (+6.0% over EATA on ViTBase), while achieving the best calibration.
-More critically, EATA-C is the only method that reduces calibration error on both ResNet-50 and ViTBase, suggesting the superiority in managing model uncertainty and data uncertainty differently for calibrated test-time adaptation.
+[…]
+Regarding calibration, existing methods consistently exhibit substantial calibration error (e.g., ROID and CoTTA are 60.46% and 27.59% on ECE with ViT-Base), suggesting miscalibration as a prevalent issue in the unsupervised test-time adaptation.
+By filtering unreliable samples to reduce noisy learning signals, EATA improves calibration over Tent (e.g., $11.7\%\rightarrow 10.5\%$ on Gaussian Noise with ResNet-50), though miscalibration is yet significant.
+By further decreasing reducible model uncertainty and reflecting data uncertainty in model predictions, our enhanced method, EATA-C, reduces the ECE of EATA by relatively 59.8% on ResNet-50 and 64.9% on ViT-Base, demonstrating the strong calibration effect of EATA-C across diverse datasets and architectures.
+In summary, while EATA-C improves performance and efficiency over the state of the art, our EATA-C further achieves high accuracy, well-calibrated prediction, and efficient computation simultaneously, establishing a new benchmark for test-time adaptation.
+[…]
 Results on CityScapes-to-ACDC.
-Following [64], we evaluate our method on semantic segmentation task in the lifelong adaptation scenario.
-From Table IV, EATA achieves comparative performance by filtering unreliable predictions and constraining the important weights from drastic change, while its counterpart, Tent, severely suffers from error accumulation.
-Moreover, by differentiating model uncertainty and data uncertainty in test-time adaptation, EATA-C achieves state-of-the-art performance, outperforming EATA by 4.6% and CoTTA by 3.2% in mIoU.
-More importantly, EATA-C can progressively accumulate knowledge over adaptation to improve its performance.
-In the first and tenth rounds of adaptation, EATA-C achieves 59.8% vs.
-62.3% in average mIoU on the four adverse datasets, further verifying the effectiveness of our method.
+Following [63], we evaluate our method on the semantic segmentation task in a lifelong adaptation scenario.
+From Table IV, while DAT initially achieves higher mIOU, it tends to overfit, leading to significant performance degradation in subsequent adaptations.
+In contrast, our EATA maintains a more stable performance compared to DAT and Tent, by filtering unreliable predictions and preventing drastic changes to important model parameters.
+Moreover, by replacing entropy minimization with our consistency maximization objective for more robust learning signals, EATA-C achieves state-of-the-art performance, surpassing EATA by 4.6% and CoTTA by 3.2% in mIoU over ten adaptation rounds.
+More critically, our EATA-C showcases consistent improvement over lifelong adaptation, increasing the average mIOU on four datasets from 59.8% (first round) to 62.3% (tenth round), further highlighting our long-term effectiveness.
 5.2 Demonstration of Preventing Forgetting
@@
 5.3 Ablation Studies
-Effect of Components in $S({\bf x})$ (Eqn.
-8).
-Our EATA accelerates test-time adaptation by excluding two types of samples out of optimization: 1) samples with high prediction entropy values (Eqn.
-3) and 2) samples that are similar (Eqn.
-8).
-We ablate both of them in Table V.
-Compared with the baseline $S({\bf x})\small{=}1$ (the same as Tent), introducing $S^{ent}({\bf x})$ in Eqn.
-(3) achieves better accuracy and fewer backwards (e.g., 49.6% (37,636) vs.
-33.4% (50,000) on level 5).
-This verifies our motivation in Figure 2 that some high-entropy samples may hurt the performance since their gradients are unreliable.
-When further removing some redundant samples that are similar (Eqn.
-8), our EATA further reduces the number of back-propagation (e.g., 37,636$\rightarrow$28,168 on level 5) and achieves comparable OOD error (e.g., 50.4% vs.
-49.6%), demonstrating the effectiveness of our sample-efficient optimization strategy.
+[…]
 Effect of Components in EATA-C.
-By further differentiating between model uncertainty and data uncertainty based on EATA, our EATA-C achieves a substantial improvement in performance, efficiency and calibration.
-We conduct ablation to verify their effectiveness in Table VI.
-Note that adapting with simply the consistency loss in Eqn.
-(12) already leads to a strong baseline.
-By re-calibrating confidence on different samples via Eqn.
-(14), we further mitigate the calibration error, e.g., 3.3% (Exp1$\rightarrow$2.5(Exp3) on ECE.
-Compared with Exp3, Exp6 maintains comparable calibration while improving accuracy (i.e., 65.9$\rightarrow$66.4%), demonstrating that our anti-forgetting regularizer does not limit OOD adaptation.
-When further introducing the sample selection in Eqn.
-(17), EATA-C is able to effectively reduce both forward passes (i.e., 50,000$\times$2$\rightarrow$82,364) and backward passes (50,000$\rightarrow$32,364) by a large margin.
-Note that although accuracy and calibration are slightly compromised in EATA-C, due to the robustness of our consistency loss on corrupted images compared with Tent (see Figure 7), its improvement over the source model (65.9% vs.39.8% on accuracy, and 3.3% vs.7.5% on ECE) is still significant.
+Our EATA-C aims to achieve a favorable balance between accuracy, calibration, and efficiency.
+We conduct an ablation study to verify the effectiveness of each module as in Table VI.
+The results indicate the following findings: 1) Consistency Loss: Incorporating the consistency loss alone substantially enhances the source model’s robustness and reduces ECE; 2) Entropy Regularization: The min-max entropy regularizer further calibrates prediction confidence and leads to a slight improvement in accuracy, e.g., accuracy increases from 48.8% (Exp 10) to 49.0% (EATA-C), and ECE decreases from 5.1% to 4.6%; 3) Fisher Regularization: This anti-forgetting regularizer contributes to TTA stability, as in the lifelong TTA of Table II and Figure 3.
+In single-domain TTA, it also positively affects both ECE and accuracy, e.g., ECE decreases from 5.4% (ETA-C) to 4.6% (EATA-C) and accuracy improves from 48.9% to 49.0%; 4) Active Sample Selection: By filtering out unreliable and redundant test samples, active sample selection significantly boosts computational efficiency while maintaining or improving accuracy, e.g., accuracy increases from 48.5% (Exp 6) to 49.0% (EATA-C) while reducing the required backward passes by 35%.
+More discussions on wall-clock time and memory usage are provided in Table VII.
+These results collectively underscore the effectiveness of each component.
 Figure 4: Effect of different entropy margins $E_{0}$ in Eqn. (3). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
+Figure 5: Effect of different similarity threshold $\epsilon$ in Eqn.( 7). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
 Figure 6: Effect of #samples for calculating Fisher in Eqn. (11). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
-Figure 4: Effect of different entropy margins $E_{0}$ in Eqn. (3). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
-Figure 5: Effect of #samples for calculating Fisher in Eqn. (11). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
-Figure 6: Effect of different stochastic depth ratios in EATA-C. Results obtained on ImageNet-C(Gaussian, level 5) with ViTBase.
+Figure 7: Effect of different balancing factor $\alpha$ in Eqn. (10). Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
+Figure 8: Effect of different balancing factor $\beta$ in Eqn. (10). Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
+Figure 9: Effect of different stochastic depth ratios in EATA-C. Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
 TABLE V: Effectiveness of components in sample-adaptive weight $S({\bf x})$ in EATA on ImageNet-C (Gaussian noise) with ResNet-50.
@@
 Baseline ($S({\bf x})\small{=}1$) 68.8 50,000 33.4 50,000
-  $+S^{ent}{({\bf x})}$ (Eqn. 3) 70.7 45,302 49.6 37,636
-  $+S{({\bf x})}$ (Eqn. 8) 70.8 36,057 50.4 28,168
-Entropy Constant $E_{0}$ in Eqn.
-(3).
-We evaluate our EATA with different $E_{0}$, selected from {0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.55}×$\times\ln 10^{3}$, where $10^{3}$ is the class number of ImageNet.
-From Figure 6, our EATA achieves excellent performance when $E_{0}$ belongs to $[0.4,0.5]$.
-Either a smaller or larger $E_{0}$ would hamper the performance.
-The reasons are mainly as follows.
-When $E_{0}$ is small, EATA removes too many samples during adaptation and thus is unable to learn enough knowledge from the remaining samples.
-When $E_{0}$ is too large, some high-entropy samples would take part but contribute unreliable and harmful gradients, resulting in performance degradation.
-As larger $E_{0}$ leads to more backward passes, we set $E_{0}$ to 0.4×$\times\ln 10^{3}$ for efficiency-performance trade-off and fix the proportion of 0.4 for all other ImageNet experiments.
+$S^{ent}{({\bf x})}$ (Eqn. 3) 70.7 45,302 49.6 37,636
+$S{({\bf x})}$ (Eqn. 8) 70.8 36,057 50.4 28,168
+[…]
 TABLE VI: Effects of components in EATA-C.
-Results obtained on 15 corruptions of ImageNet-C (severity level 5) with VitBase in single-domain adaptation scenario, i.e., the model parameters are reset before adapting to a new corruption.
+Results obtained on 15 corruptions of ImageNet-C (level 5) with ResNet-50 in single-domain TTA scenario, i.e., the model parameters are reset before adapting to a new domain.
 CL denotes consistency loss.
-ER denotes min-max entropy regularization.
-FR denotes anti-forgetting Fisher regularization.
-SS denotes active sample selection.
-Experiment CL ER FR SS Average
+ER denotes min-max entropy regularizer.
+FR denotes Fisher regularizer.
+$S$ denotes active sample selection.
+Experiment CL ER FR $S$ Average
 Acc. ECE #Forwards #Backwards
-1 65.6 3.3 50,000$\times$2 50,000
-2 53.4 4.1 50,000$\times$2 50,000
-3 65.9 2.5 50,000$\times$2 50,000
-4 66.0 3.3 50,000$\times$2 50,000
-5 53.0 4.5 50,000$\times$2 50,000
-6 66.4 2.6 50,000$\times$2 50,000
-7 65.6 4.4 83,182 33,182
-8 52.9 3.4 73,323 23,323
-9 (ETA-C) 65.3 3.4 82,015 32,015
-10 66.2 4.3 83,429 33,429
-11 53.2 3.4 73,723 23,723
-12 (EATA-C) 65.9 3.3 82,364 32,364
+Source 39.8 7.5 50,000 0
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+7 48.9 5.9 83,583 33,583
+8 42.9 4.4 74,705 24,705
+9 (ETA-C) 48.9 5.4 82,882 32,882
+10 48.8 5.1 82,952 32,952
+11 42.3 3.8 74,226 24,226
+12 (EATA-C) 49.0 4.6 82,492 32,492
+Similarity Threshold $\epsilon$ in Eqn.
+7.
+We use $\epsilon$ to select diverse samples for TTA.
+From Figure 6, EATA maintains stable accuracy across a wide range of $\epsilon\in[0.01,0.08]$, showcasing insensitivity, while a smaller $\epsilon$ removes significantly more samples and improves computational efficiency.
+We set $\epsilon{=}0.05$ without careful tuning.
+More results on efficiency (i.e., time and memory usage) of EATA and EATA-C with varying $\epsilon$ are provided in Table VII.
 Number of Samples for Calculating Fisher in Eqn.
@@
 These results show that our EATA does not need to collect too many ID samples, which are easy to obtain in practice.
+Factor $\alpha$ for Entropy Regularizer in Eqn.
+16.
+We directly set $\alpha{=}0.1$ to align the magnitudes of consistency loss and entropy regularization loss for EATA-C without careful tuning.
+From Figure 9, increasing $\alpha$ within $[0,0.1]$ effectively reduces more ECE while maintaining stable accuracy, verifying its efficacy.
+However, when $\alpha$ exceeds 0.1, the entropy regularization loss dominates the adaptation, which leads to a gradual decline in accuracy.
+Factor $\beta$ for Fisher Regularizer in Eqn.
+16.
+From Figure 9, compared to ETA-C (i.e., setting $\beta{=}0$), introducing the fisher regularizer consistently achieves better accuracy.
+Moreover, once activated, the performance of EATA becomes largely insensitive to $\beta$ within the tested range of $[20,140]$, highlighting its robustness.
 Stochastic Depth Ratio for Obtaining Sub-Network.
@@
 5.4 More Discussions
-Results on Single Sample Adaptation (Batch Size $B=1$).
-In our proposed methods, we conduct test-time learning over a batch of test samples each time.
-When batch size $B$ equals 1, directly applying Tent and our ETA on models with batch normalization layers may fail (see Table VII), but it is not a very significant issue in practice.
-Actually, one can address this by either maintaining a sliding window $\{{\bf x}_{i}\}_{i=t-L}^{t}$ including $L$ previous samples and the current test sample ${\bf x}_{t}$, or data augmentation techniques adopted in existing single sample TTA methods ([14, 33]).
-The first way is simple to implement and incurs only minor extra costs.
-From Table VII, our ETA with sliding window (ETA-wnd) works well and consistently outperforms the Tent counterpart.
-Wall-clock Time Speed-up of EATA and EATA-C.
-In the current PyTorch version, gradient computation is conducted on the full mini-batch, even if instance-wise masks are applied.
-To benefit from active sample selection, in EATA, we perform one forward-only pass with $B$ samples, and one extra forward and backward pass with $B^{\prime}$ selected samples.
-We achieve wall-time speed-up when setting $E_{0}$ in Eqn.
-(3) to $0.1\times\ln C$, where $C$ is the number of task classes.
-For ViTBase on ImageNet-C (Gaussian noise, level=5) with one Titan Xp GPU, the actual run time is 518s for Tent (33.3% accuracy), 346s for EATA (39.5% accuracy), and 402s for EATA-C (51.3% accuracy).
-It’s also worth mentioning that in EATA-C, the sub-network’s computation is not accelerated even if we select partial samples for each layer due to a similar setback.
-Actually, this is a more engineering problem, and an ideal implementation should further speed up both EATA and EATA-C.
+Efficiency Analysis of EATA and EATA-C.
+We evaluate the efficiency of our methods by including a more comprehensive comparison of time and memory usage for TTA, as in Table VII.
+The results reveal the following: 1) Adaptation Time: EATA and EATA-C require significantly less adaptation time than most baseline methods, while maintaining competitive or superior accuracy.
+For example, on ViT-Base, EATA-C improves accuracy from 48.8% (ROID) to 56.8% with a reduction in adaptation time from 181.6 seconds to 114.9 seconds.
+Moreover, by setting a stricter threshold $\epsilon$ to filter redundant test samples, EATA and EATA-C can be further accelerated while maintaining performance.
+For example, using $\epsilon{=}0.02$, EATA on ViT-Base increases accuracy from 33.4% (Tent) to 50.7% while reducing adaptation time from 106.1 seconds to 85.1 seconds; 2) Memory Usage: Our methods also demonstrate efficient memory utilization, where EATA and EATA-C consume substantially less memory compared to all competing TTA methods, e.g., on ResNet-50, memory usage decreases from 5417.7MB (SAR) to 2205.5MB (EATA-C, $\epsilon{=}0.02$) while accuracy increases from 29.6% to 35.9%.
+This memory reduction is achieved through our active sample selection strategy, which reduces the number of samples involved in backpropagation.
+Note that the current Pytorch implementation does not support instance-wise gradient computation, thus an ideal implementation should further speed up both EATA and EATA-C.
+See more discussions on our implementation details in Appendix A.
+TABLE VII: Efficiency comparison regarding wall-clock time and peak memory usage on ImageNet-C (Gaussian, severity level 5) with an A100 GPU.
+Method ResNet-50 ViT-Base
+Acc. (%) Time (s) Mem. (MB) Acc. (%) Time (s) Mem. (MB)
+Source 1.8 54.6 771.7 12.9 55.7 816.6
+Tent [12] 28.0 99.9 5417.6 33.4 106.1 7433.2
+SAR [18] 29.6 149.0 5417.7 43.1 167.3 7433.2
+ROID [64] 36.7 190.2 9315.3 48.8 181.6 12321.4
+CoTTA [16] 19.9 241.0 12196.7 45.6 229.9 22296.5
+TEA [65] 18.5 2266.1 15942.0 46.9 - -
+MEMO [14] 6.8 36329.2 8154.3 32.2 39918.8 11061.1
+EATA ($\epsilon{=}0.02$) 35.2 93.0 2285.2 50.7 85.1 4234.6
+EATA-C ($\epsilon{=}0.02$) 35.9 95.2 2205.5 55.0 88.8 4014.4
+EATA (Ours) 35.6 106.5 3693.0 50.5 108.7 5887.6
+EATA-C (Ours) 37.2 122.4 3358.7 56.8 114.9 5786.6
+TABLE VIII: Reliability of data uncertainty indicator.
+We report sub-model Acc.
+(%) on ImageNet-C(Gaussian, level 5) after adapting to $B$ batches (batch size 64) with ViT-Base.
+“#Uncertain” are samples with disagreed predictions.
+“Indicator Acc.” is the ratio of these samples misclassified by sub-model.
+Metric
+Source
+EATA-C
+EATA-C
+EATA-C
+EATA-C
+EATA-C
+(B=150B{=}150) (B=300B{=}300) (B=450B{=}450) (B=600B{=}600) (B=750B{=}750)
+Model Acc. (%) 9.3 47.0 50.2 50.7 51.9 52.0
+#Uncertain 26144 15697 14682 14149 13648 13722
+Indicator Acc. (%) 97.1 90.0 88.9 89.3 88.6 89.1
+Figure 10: Calibration comparison of each confidence interval on ImageNet-C(Fog, level 5) with ViT-Base. Results are visualized following [44].
+Effectiveness of Data Uncertainty Indicator.
+We evaluate the effectiveness of our data uncertainty indicator, i.e., Eqn.
+(15), throughout TTA.
+The results are detailed in Table VIII: 1) Consistently High Indicator Accuracy: Across various adaptation stages, our indicator continues to reliably identify uncertain samples, on which the sub-model’s prediction is likely to be incorrect.
+Specifically, the indicator maintains around 90% accuracy, suggesting its effectiveness even in the early stage of adaptation.
+This reliability allows us to apply entropy maximization on these uncertain data to improve calibration without hindering adaptation; 2) Reduced Uncertain Samples Over Time: The model’s initial poor performance mainly leads to a higher number of uncertain samples, including inherently difficult data for discrimination and data the model has yet to learn.
+However, model uncertainty quickly explains away during TTA (i.e., within one-fifth of the data stream), leading to a stabilized number of uncertain samples that reflects irreducible data uncertainty.
 Additional Memory by Fisher Regularizer.
@@
 We evaluate Tent and our EATA/EATA-C on mixed ImageNet-C (level=3 or 5) that consists of 15 different corruption types/distribution shifts (totaling 750k images).
-Results in Table VIII show the stability of EATA and EATA-C in large-scale and complex TTA scenarios.
-TABLE VII: Effectiveness of ETA under sliding window strategy (with different window length $L$) for single sample adaptation with ResNet-50.
-We report corruption accuracy (%) on ImageNet-C (Gaussian noise, level 5).
-Base ResNet-50 Tent ETA Tent-wnd ($L\small{=}32)$ ETA-wnd ($L\small{=}32$) Tent-wnd ($L\small{=}64$) ETA-wnd ($L\small{=}64$)
-2.2 0.1 0.1 28.1 30.8 29.5 32.4
-TABLE VIII: Comparison with Tent [12] w.r.t. corruption accuracy (%) with mixture of 15 corruption types on ImageNet-C with ViTBase.
- Severity Source Tent EATA (ours) EATA-C (ours)
-Level=3 63.0 69.8${}_{(+6.8)}$ 71.1${}_{(+8.1)}$ 72.4${}_{(+9.4)}$
-Level=5 39.8 47.0${}_{(+7.2)}$ 58.2${}_{(+18.4)}$ 60.4${}_{(+20.6)}$
+Results in Table IX show the stability of EATA and EATA-C in large-scale and complex TTA scenarios.
+TABLE IX: Comparison with Tent [12] w.r.t. corruption accuracy (%) with mixture of 15 corruption types on ImageNet-C with ViT-Base.
+ Severity
+Source
+Tent
+EATA (ours)
+EATA-C (ours)
+Level=3 63.0 69.8(+6.8) 71.1(+8.1) 72.4(+9.4)
+Level=5 39.8 47.0(+7.2) 58.2(+18.4) 60.4(+20.6)
 (a) Performance and calibration throughout adaptation
@@
 Figure 11: Comparisons of performance, calibration and stability using consistency loss in Eqn.
-(12) or entropy minimization loss (i.e., Tent) on ImageNet-C (Gaussian noise, severity level=5) with ViTBase.
+(12) or entropy minimization loss (i.e., Tent) on ImageNet-C (Gaussian noise, severity level=5) with ViT-Base.
 In Left, we split the dataset into 80% and 20% slices to conduct adaptation and to evaluate the adapted model.
-We report the performance and the calibration of the adapted model every 75 batches.
-In the Middle and Right, we evaluate the stability of using consistency loss and Tent under various combinations of learning rates and adaptation steps per batch.
+We report the performance and the calibration of the adapted model every 75 batches with a batch size of 64.
+In the Middle and Right, we evaluate the stability of consistency loss and Tent under various combinations of learning rates and adaptation steps per batch.
 Consistency loss achieves substantially higher OOD performance and better stability while maintaining lower ECE.
-Advantage of Consistency Loss over Tent [12].
-We conduct a comprehensive comparison of the performance, calibration and stability between the use of consistency loss, as defined in Eqn.
-(12), and entropy minimization loss (i.e., Tent [12]) to reduce uncertainty at test time.
-The experiments are conducted on ImageNet-C (Gaussian noise, level 5) with ViTBase.
-In Figure 7(a), we partition the dataset into 80% and 20% for adaptation and evaluation, respectively, to show how adaptation impacts performance and calibration over time.
-Analyzing Figure 7(a), we discern several key findings: 1) consistency loss consistently demonstrates superior performance and calibration throughout adaptation.
+Calibration Across Confidence Intervals.
+EATA-C aims to achieve a favorable balance among accuracy, calibration, and efficiency.
+In EATA-C, we discard high-entropy samples (termed active sample selection) mainly to improve computational efficiency.
+While some high-entropy samples might benefit from further calibration, they typically yield unreliable pseudo-labels, which can negatively impact the stability and effectiveness of TTA (see Table VI).
+Instead of directly calibrating on these samples, we show that focusing adaptation and calibration on only low-entropy samples can also improve the calibration on high-entropy ones, as in Figure 10, while improving TTA efficiency and effectiveness.
+Advantage of Consistency Maximization over Tent [12].
+We conduct a comprehensive comparison w.r.t.
+performance, calibration, and stability between the use of consistency loss, as defined in Eqn.
+(12), and Tent [12] to reduce uncertainty during testing.
+From Figure 11(a), we have the following observations: 1) Consistency loss consistently demonstrates superior performance and calibration throughout adaptation.
 2) Consistency loss is more sample-efficient, where adapting with as few as 75 batches can significantly outperform Tent [12] that adapts with 300 batches.
-3) Tent [12] shows rapid degradation in performance and calibration after convergence, which indicates that it tends to over-fit.
-In contrast, consistency loss maintains stable performance and calibration after convergence and continuously exhibits strong generalization throughout adaptation.
-In Figures 7(b) and 7(c), we study the stability of consistency loss and Tent [12] across various combinations of learning rates and adaptation steps following [71].
-As illustrated in Figure 7(b), consistency loss demonstrates remarkable performance stability and benefits from the increment of adaptation steps (e.g., $55.8\%\rightarrow 58.6\%$, the best performance under 1 and 5 adaptation steps, respectively).
-In contrast, Tent [12] is sensitive to the combination of learning rate and adaptation steps, where its performance may deteriorate to as low as 1%.
-Such a phenomenon further highlights Tent [12]’s tendency to over-fit.
-In this sense, consistency loss emerges as the superior choice considering performance, calibration, stability and sample efficiency.
+3) Tent [12] shows rapid degradation in performance and calibration after convergence.
+In contrast, consistency loss maintains stable performance and calibration after convergence and continuously exhibits strong generalization throughout TTA.
+We further evaluate the stability of consistency loss and Tent [12] across various combinations of learning rates and adaptation steps in Figures 11(b) and 11(c) following [73].
+The results highlight that our consistency loss demonstrates remarkable stability and benefits from increased adaptation steps (e.g., $55.8\%\rightarrow 58.6\%$, the best performance under 1 and 5 adaptation steps, respectively).
+In contrast, Tent [12] is highly sensitive to the combination of learning rate and adaptation steps, where its performance may deteriorate to as low as 1%, further indicating its tendency to overfit.
+These findings collectively underscore the superiority of our consistency loss regarding performance, calibration, and stability, making it a more robust choice for TTA.
 6 Conclusion
@@
 Extensive experimental results on image classification and semantic segmentation demonstrate the effectiveness of our proposed methods.
+Acknowledgments
+This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), and TCL Science and Technology Innovation Fund, China.
 Supplementary Materials for “Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting”
@@
   • In Section A, we provide more details of our proposed EATA and EATA-C.
-  • In Section B, we show more experimental results to compare the out-of-distribution performance, calibration and efficiency with state-of-the-art methods on ImageNet-C with different corruption types.
-  • In Section C, we give more experimental results to demonstrate the anti-forgetting ability of our EATA.
-  • In Section D, we provide more discussions on related training-time robustification studies.
+  • In Section B, we present additional experimental results and ablation studies to further demonstrate our superiority in out-of-distribution performance, calibration, efficiency, and stability.
+  • In Section C, we provide further insights, and empirical/visual evidence to validate the design choices in EATA-C.
 Appendix A More Details of EATA and EATA-C
@@
 A.2 More Details on Datasets
-Following the settings of Tent [12] and MEMO [14], we conduct experiments on four benchmark datasets for out-of-distribution generalization, i.e., CIFAR-10-C, ImageNet-C [9], ImageNet-R [65], and ACDC [66].
-CIFAR-10-C and ImageNet-C consist of corrupted versions of the validation images on CIFAR-10 [72] and ImageNet [73], respectively.
-The corruptions include 15 diverse types of 4 main categories (i.e., noise, blur, weather, and digital).
-Each corruption type has 5 different levels of severity.
-ImageNet-R contains 30,000 images with various artistic renditions of 200 ImageNet classes, which are primarily collected from Flickr and filtered by Amazon MTurk annotators.
-ACDC contains four categories of images collected in adverse conditions, including fog, night, rain, and snow. Following [64], we use 400 unlabeled images from each adverse condition for adaptation.
-A.3 More Experimental Protocols
-Our EATA and EATA-C.
-Following Tent [12] and CoTTA [64], we use ResNet-26 for CIFAR-10 experiments, ResNet-50 and ViTBase for ImageNet experiments, and Segformer-B5 for ACDC experiments.
-In classification experiments, the models are trained on the original CIFAR-10 or ImageNet training set and then tested on clean or the aforementioned OOD test sets.
+Following the settings of Tent [12] and MEMO [14], we conduct experiments on three benchmark datasets for out-of-distribution generalization, i.e., ImageNet-C [9], ImageNet-R [66], and ACDC [67].
+ImageNet-C consists of various versions of corruption applied to 50,000 validation images from ImageNet.
+The dataset encompasses 15 distinct corruption types of 4 main categories, including Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness, contrast, elastic transformation, pixelation, and JPEG compression.
+Each corruption type is characterized by 5 different levels of severity, with higher severity levels indicating a more severe distribution shift.
+ImageNet-R contains 30,000 images with various artistic renditions of 200 ImageNet classes. These images are primarily collected from Flickr and filtered by Amazon MTurk annotators.
+ACDC contains four categories of images collected in adverse conditions, including fog, night, rain, and snow.
+Following CoTTA [63], we use 400 unlabeled images from each adverse condition for continuous test-time adaptation.
+A.3 More Experimental Protocols on Evaluation
+Following Tent [12] and CoTTA [63], we use ResNet-50 and ViT-Base for ImageNet experiments, and Segformer-B5 for ACDC experiments.
+In classification experiments, the models are trained on the original ImageNet training set and then tested on clean or the aforementioned OOD test sets.
 In semantic segmentation experiments, the model is trained on Cityscapes and then tested on ACDC.
-For a fair comparison, the parameters of ViTBase and Segformer-B5 are directly obtained from the timm [^1] https://github.com/pprp/timm and Segformer [68], respectively.
+For a fair comparison, the parameters of ViTBase and Segformer-B5 are directly obtained from the timm [^1] https://github.com/pprp/timm and Segformer [^2] https://github.com/NVlabs/SegFormer [68], respectively.
 ResNet-50 is trained via the official code in the torchvision [^3] https://github.com/pytorch/vision library with stochastic depth.
-ResNet-26 is trained via the official code of TTT by the same hyper-parameters, replacing the group norm with the batch norm, and removing the rotation head.
-For test time adaptation, we use SGD as the update rule, with a momentum of 0.9 and a batch size of 64.
-In EATA and ETA, the learning rate is set to 0.005 on CIFAR-10, 0.00025/0.001 for ResNet-50/ViTBase on ImageNet, and 7.5×$\times 10^{-5}$ on ACDC, respectively (following Tent, SAR and CoTTA).
-The entropy constant $E_{0}$ in Eqn.
-(3) is set to $0.4\times\ln C$ in classification experiments and $0.1\times\ln C$ in segmentation, where $C$ is the number of task classes.
-The $\epsilon$ in Eqn.
-(8) is set to 0.4 for CIFAR-10, and 0.05 for ImageNet and ACDC.
-The trade-off parameter $\beta$ in Eqn.
-(10) is set to 1/2,000/500 for CIFAR-10/ImageNet/ACDC to make two losses have a similar magnitude.
-In EATA-C, the learning rate is set to 0.005/0.1 for ResNet-50/ViTBase on ImageNet, and 0.0005 on ACDC, respectively.
-The sub-network is obtained via stochastic depth regularization [63] with a drop ratio of 0.2/0.6 for ImageNet/ACDC.
-The entropy constant $E_{0}$ in Eqn.
-(3) is set to 0.4/0.5×$\times\ln C$, and the $\epsilon$ in Eqn.
-(8) is set to 0.05/0.07 for ViTBase/ResNet-50 on ImageNet.
-In Eqn.
-(16), the trade-off parameter $\alpha$ is set to 50/500 on ImageNet/ACDC and $\beta$ is set to 0.1 on ImageNet.
-Note that for the semantic segmentation task, our EATA-C does not perform min-max entropy regularization and active sample selection due to the long-tailed problem [74, 75].
-For both EATA and EATA-C, the moving average factor $\alpha$ in Eqn.
-(6) is set to 0.1, and we use 2,000/20 samples for ImageNet/ACDC to calculate $\omega(\theta_{i})$ in Eqn.
-(11).
+[…]
 Compared Methods.
-For TTA [76], BN adaptation [32], MEMO [14], CoTTA [64], and SAR [18] the hyper-parameters follow their original papers or MEMO.
-Specifically, the augmentation size of TTA [76] is set to 32 and 64 for CIFAR-10 and ImageNet, respectively.
-For BN adaptation [32], both the batch size $B$ and prior strength $N$ are set to 64.
+For BN adaptation [29], MEMO [14], CoTTA [63], SAR [18], ROID [64], Rdump [71], TEA [65], and DAT [70], the hyper-parameters follow their original papers or MEMO.
+Specifically, for BN adaptation [29], both the batch size $B$ and prior strength $N$ are set to 64.
 The learning rate in CoTTA [63] for ViTBase is set to 0.005, and the augmentation threshold $p_{th}$ is set to 0.1.
-Other hyper-parameter settings of CoTTA, and the hyper-parameter settings of MEMO [14] and SAR [18] can be found in their original paper.
-For Tent [12], we use SGD as the update rule with a momentum of 0.9.
-The batch size is 64 for both ImageNet and CIFAR-10 experiments.
-The learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase on ImageNet and 0.005 on CIFAR-10, respectively.
-Note that the hyper-parameters of Tent are totally the same as our EATA for a fair comparison.
-For TTT [11], we strictly follow their original settings except for the augmentation size at test time for ImageNet experiments.
-According to TTT’s implementation, the augmentation size is set to 64, which, however, is very time-consuming (e.g., about 12 GPU hours on a single Tesla V100 GPU on ImageNet-C with a specific corruption type and severity level).
-In our implementation, we decrease this augmentation size to 20, which has only a slight performance difference compared with augmentation size 64.
-For example, the performances of TTT on ImageNet-C (Gaussian noise, severity level 5) with ResNet-18 are 26.2% (64) vs.
-26.0% (20).
-We recommend the original papers of the above methods to readers for more implementation details.
-Appendix B More Results on Out-of-distribution Performance, Efficiency and Calibration
-In Table D, we provide more results to compare our EATA and EATA-C with state-of-the-art methods in the single-domain adaptation scenario on ImageNet-C with the severity level 5.
-Although the compared methods exhibit stronger performance in the less challenging adaptation scenario, our EATA and EATA-C still demonstrate substantial superiority.
-Specifically, by filtering the unreliable and redundant samples from adaptation, our EATA achieves significant accuracy improvement while requiring less backward-propagation (e.g., 55.5% (72,446) vs.61.8% (32,524) over the best method SAR on ViTBase).
-Meanwhile, EATA-C further enhances performance while mitigating calibration error on both ResNet-50 and ViTBase, verifying the promising effectiveness of our EATA and EATA-C across different adaptation scenarios.
-In Table E, we provide the complete results for the lifelong test-time adaptation on ACDC.
-Compared with Tent, our EATA effectively alleviates error accumulation by excluding unreliable gradients and constraining important weights from drastic changes.
-Furthermore, besides maintaining strong performance in the long term, our EATA-C is able to further accumulate knowledge to improve performance over adaptation, demonstrating strong potential.
-In Figure B, we show the number of backward propagation of our ETA on ImageNet-C with different corruption types and severity levels.
-Across various corruption types, our ETA shows great superiority over existing methods in terms of adaptation efficiency.
-Compared with MEMO (50,000$\times$64) and Tent (50,000), our ETA only requires 31,741 backward passes (averaged over 15 corruption types) when the severity level is set to 3.
-The reason is that we exclude some unreliable and redundant test samples out of test-time optimization.
-In this case, we only need to perform backward computation on those remaining test samples, leading to improved efficiency.
-In Table F, we compared ETA and EATA with existing methods on CIFAR-10-C.
-Under the same base model (ResNet-26 with batch normalization), ETA achieves a lower average error than Tent (19.4% vs.
-20.2%) with fewer requirements of back-propagation (8,192 vs.
-10,000).
-Moreover, the performance gain over the base model of ETA is larger than that of TTT, i.e., +9.0% vs.
-+7.2% on accuracy.
-These results are consistent with the ones on ImageNet-C/R, further demonstrating the effectiveness and superiority of our method.
-Comparison with Tent using Different Learning Rates.
-In our sample-adaptive weight $S{({\bf x})}$ in Eqn.
-(8), each test sample has a specific weight $S{({\bf x})}$ and the value of $S{({\bf x})}$ is always larger than 1.
-In this sense, training with sample-adaptive weight $S{({\bf x})}$ indeed has the same effect as training with larger learning rates.
-Therefore, we compare our EATA with the baseline (Tent) using different learning rates.
-We increase the learning rate from $2.5\times 10^{-4}$ (which is the default of Tent) to $25.0\times 10^{-4}$ and report results in Table A.
-With the learning rate increasing from $2.5\times 10^{-4}$ to $10.0\times 10^{-4}$, the error of Tent decreases from 45.3% to 43.9%, indicating that a larger learning rate may enhance the performance in some cases.
-However, when the learning rate becomes larger to $20.0\times 10^{-4}$, the performance of Tent degrades.
-More critically, our EATA method outperforms Tent with varying learning rates.
-These results verify that simply enlarging the learning rate is not able to achieve competitive performance with our proposed sample-adaptive adaptation method, demonstrating our superiority.
-TABLE A: Comparison with Tent under different learning rates (×$\times 10^{-4}$) on ImageNet-C (Gaussian noise) regarding Error (%).
-EATA (ours) Tent [12]
-Severity $lr=2.5$ $lr=2.5$ $lr=10.0$ $lr=20.0$ $lr=25.0$
-Level 3 42.6 45.3 43.9 44.4 45.1
-Level 5 65.0 71.6 72.2 83.6 87.1
-EATA under Different Random Orders.
-Table B records EATA’s performance (mean & stdev.) on randomly shuffled test samples with 10 different random seeds (from 2020 to 2029).
-From the results, EATA performs consistently across different random orders, showing the stability of EATA.
-TABLE B: The mean and stdev. of corruption accuracy (%) of EATA over 10 random orders, on ImageNet-C (level 5) with ResNet-50.
-Gauss. Shot Impul. Defoc. Glass Motion Zoom Avg.
-34.9±${}_{\pm 0.2}$ 36.9±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$ 33.6±${}_{\pm 0.3}$ 33.3±${}_{\pm 0.2}$ 47.2±${}_{\pm 0.3}$ 52.7±${}_{\pm 0.1}$ 35.8±${}_{\pm 0.2}$
-EATA with Different Numbers of Test Samples.
-We investigate the effect of the total number of test samples ($N$) in EATA, where fewer samples are sampled from the entire test set.
-From Table C, EATA works well and consistently outperforms Tent counterpart, regardless of the number of test samples.
-Meanwhile, if there are many test samples, EATA would benefit more, i.e., larger accuracy gain.
-TABLE C: Comparison with Tent [12] w.r.t. corruption accuracy (%) with fewer ($N$) test samples on ImageNet-C (Gaussian noise, level 5) with ResNet-50.
-Method $N\small{=}$256 $N\small{=}$512 $N\small{=}$1,024 $N\small{=}$2,048 $N\small{=}$4,096 $N\small{=}$10,000
-Tent 13.7 18.2 14.8 15.2 17.5 20.9
-EATA 14.8 20.1 16.3 18.7 23.5 27.6
-Appendix C More Results on Prevent Forgetting
-In this section, we provide more results to demonstrate the effectiveness of our EATA in preventing forgetting.
+TEA [65] uses SGD with a learning rate of 0.00025/0.005 for ResNet-50/ViT-Base, respectively.
+Other hyper-parameter settings of CoTTA and TEA, as well as those for MEMO, SAR, ROID, Rdump, and DAT, can be found in their original paper.
+For Tent [12], we use SGD with a momentum of 0.9.
+The batch size is 64 for both ImageNet experiments and the learning rate is set to 0.00025/0.001 for ResNet-50/ViTBase, respectively.
+Note that the hyper-parameters of Tent are the same as our EATA for a fair comparison.
+Appendix B More Experimental Results
+B.1 Results under Single-Domain Adaptation
+In Table A, we provide additional results to evaluate the effectiveness of our EATA and EATA-C under single-domain adaptation scenarios.
+From the results, while the compared methods exhibit stronger performance in the less challenging adaptation setting, our EATA and EATA-C still demonstrate substantial superiority.
+Specifically, by filtering the unreliable and redundant samples from adaptation, our EATA achieves substantial accuracy improvement while requiring significantly fewer backpropagations, e.g., on ViT-Base, average accuracy increases from 55.5% (SAR) to 61.8% (EATA) with a reduction in backward passes from 72,446 to 32,524.
+Meanwhile, EATA-C further enhances performance and reduces calibration error on both ResNet-50 and ViTBase.
+These results are consistent with our findings in lifelong adaptation (Table II), verifying the effectiveness of our EATA and EATA-C across diverse adaptation scenarios.
+B.2 More Results on Semantic Segmentation
+In Table B, we present the complete results for lifelong test-time adaptation on the ACDC dataset.
+Although lifelong adaptation offers exposure to a broader range of data with the potential for improved performance, Tent and DAT tend to suffer from error accumulation and catastrophic forgetting, leading to a decline in subsequent adaptations.
+In contrast, our EATA mitigates these issues by excluding unreliable gradients and constraining important weights from drastic changes, thereby maintaining a more stable performance.
+Moreover, our EATA-C is able to accumulate learned knowledge from prior adaptation and showcase consistent improvement over lifelong adaptation, increasing the average mIOU on four adverse datasets from 59.8% (first round) to 62.3% (tenth round).
+These results further highlight the long-term effectiveness of our methods in leveraging the potential of lifelong test-time adaptation.
+B.3 Results under Diverse Severity Levels
+[…]
+B.4 More Results on Prevent Forgetting
+We provide more results to demonstrate the effectiveness of our EATA in preventing forgetting.
 We report the comparison results of EATA (lifelong) vs.
 Tent (lifelong) and EATA vs.
-Tent in Figures C and D, respectively.
-In the lifelong adaptation scenario, Tent suffers more severe ID performance degradation than that of reset adaptation (i.e., Figure D), showing that the more optimization steps, the more severe forgetting.
+Tent in Figures F and G, respectively.
+In the lifelong adaptation scenario, Tent suffers more severe ID performance degradation than that of reset adaptation (i.e., Figure G), showing that the more optimization steps, the more severe forgetting.
 Moreover, with the increase of the severity level, the ID clean accuracy degradation of Tent increases accordingly.
@@
 In contrast, our methods achieve higher OOD corruption accuracy and, meanwhile maintain the ID clean accuracy (competitive to the original accuracy that is tested before any OOD adaptation) in both two adaptation scenarios (reset and lifelong).
-These results are consistent with that in the main paper and further demonstrate the effectiveness of our proposed anti-forgetting weight regularization.
-Appendix D More Discussions on Related Training-Time Robustification
-To defend against distribution shifts, many prior studies seek to enlarge the training data distribution to enable it to cover the possible shift that might be encountered at test time, such as adversarial training strategies [77, 78, 79], various data augmentation techniques [80, 81, 82, 65, 83] and searching/enhancing sub-networks of a deep model [84, 85].
-However, it is hard to anticipate all possible test shifts at training time.
-In contrast, we seek to conquer this test distribution shift by directly learning from test data.
-TABLE D: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$).
+These results are consistent with those in the main paper and further demonstrate the effectiveness of our proposed anti-forgetting weight regularization.
+B.5 More Ablation Studies on EATA-C’s Components
+EATA-C aims to achieve a favorable balance among accuracy, calibration, and efficiency.
+We supplement more ablation studies on different datasets and architectures to verify the effectiveness of each component, as in Table C.
+Our findings on ImageNet-R with ResNet-50 are consistent with those in Table VI, including 1) Using the consistency loss alone already yields a strong baseline; 2) Adding the entropy regularizer and Fisher regularizer further improves TTA accuracy while reducing ECE; 3) By filtering out unreliable and redundant samples, active sample selection significantly boosts computational efficiency while maintaining or improving accuracy.
+Moreover, we provide further explanations regarding the performance variations between ViT-Base and ResNet-50 when incorporating active sample selection.
+This is primarily attributed to the quality of pseudo labels for TTA.
+ViT-Base generally yields higher-quality pseudo labels, with consistently higher accuracy and lower ECE.
+Considering adaptation efficiency, the active sample selection strategy may remove a few instances that could have been beneficial to adaptation, resulting in a small drop in accuracy (e.g., from 66.4% to 65.9%), while reducing the required backward passes by over 35%.
+In contrast, ResNet-50 tends to produce less stable pseudo labels, which could adversely affect the performance if all samples are used indiscriminately, as evidenced by the collapse of TEA [65] on ResNet in Table II.
+In this case, active sample selection not only reduces unnecessary backward passes, but also enhances accuracy by focusing test-time adaptation on more trustworthy samples.
+Appendix C Additional Discussions
+C.1 Foundations of Disagreement-Based Data Uncertainty Indicator
+In EATA-C, we devise a disagreement-based indicator to identify uncertain samples.
+This is inspired by margin-based theories in machine learning [61, 62], which indicate that samples near decision boundaries are inherently more uncertain and have been well justified with theoretical guarantees.
+To be specific, we leverage prediction disagreement between the full network and sub-network to detect these uncertain boundary samples, where conflicting predictions are more likely to occur.
+Figure C provides a toy illustration of this principle and we depict our model and data uncertainty based on Figure C below.
+1) Model uncertainty.
+In our method, activating only a subset of parameters results in a new decision boundary for the sub-model, e.g., Sub-Model 1/2.
+Our consistency loss aligns the sub-model boundary with that of the full model, thereby aiming to minimize the model uncertainty and obtain an optimal decision boundary.
+2) Data uncertainty.
+Unlike our consistency loss that measures model uncertainty, in which samples across the entire data space may yield low consistency loss, data uncertainty is reflected more prominently through the disagreement of predictions near the decision boundary.
+Moreover, we provide empirical evidence of using prediction disagreement as the data uncertainty indicator in Table VIII and Figure C.
+From the results, we reveal that a substantial number of samples (over 13,000) exhibit persistent prediction disagreement even after extensive adaptation.
+Notably, about 90% of these samples are misclassified by the well-trained model, suggesting intrinsic noise or ambiguity near decision boundaries, aligned with the definition of data uncertainty.
+Collectively, the margin-based theories and our empirical results provide a sound basis for adopting disagreement as a practical indicator of data uncertainty in TTA.
+Figure B: Toy illustration. Uncertain data near the decision boundary tend to cause disagreed predictions among the sub- and full model.
+Figure C: Evolution of uncertain samples (i.e., with disagreed predictions) during adaptation on ImageNet-C (Gauss, level 5) using Eqn. (10).
+C.2 Parameter Update Strategies: Norm-Only vs. Full Updates
+In our main experiments, EATA-C updates only the normalization layers by following established practices of existing TTA work [12, 65].
+Here, we further validate this design choice by comparing it against a variant that updates all model parameters, namely EATA-C .
+The results in Table D highlight the following.
+1) Reduced Network Divergence: Updating the normalization layer effectively decreases the divergence loss between the full network and the sub-network, from 0.565 (source model) to 0.451; 2) Improved Accuracy and Calibration: While updating all model parameters yields a marginally lower divergence, our EATA-C that updates only the norm layers leads to higher accuracy (56.8% vs.
+56.0%) and lower ECE (5.2% vs.
+5.9%).
+This improvement is mainly attributed to the increased stability of TTA, where restricting updates to the normalization layers acts as a regularizer for online adaptation and reduces the risk of forgetting and overfitting by keeping the majority of parameters frozen, also known as parameter-efficient tuning.
+3) Lower Overhead: Our EATA-C also demonstrates reduced adaptation time (114.9s vs.
+140.9s) and memory usage (5786.6MB vs.
+9920.5MB) compared to EATA-C, making it more practical for real-world applications.
+C.3 Visualization of Reliable and Unreliable Test Samples
+We show some selected and removed test samples for TTA according to our entropy-based indicator in Eqn.
+(3).
+From Figure E, the selected samples generally exhibit higher image quality (i.e., with clear and distinguishable class features), where the model can generate more reliable pseudo labels for adaptation.
+In contrast, the removed samples are often ambiguous even for human interpretation, leading to potentially incorrect predictions that would introduce noisy learning signals when used for adaptation.
+This highlights the effectiveness of our entropy-based identification in improving the robustness of TTA.
+TABLE A: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, $\uparrow$) and Expected Calibration Error($\downarrow$).
 “BN” and “LN” denote batch and layer normalization.
@@
 Noise Blur Weather Digital Average
-Model Method Metric Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel JPEG Avg. # Forwards # Backwards
+Model Method Metric Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel JPEG Avg. #Forwards #Backards
 R-50 (BN) Source Acc. 1.8 3.0 1.7 18.2 10.1 13.4 20.8 14.0 22.1 21.9 58.7 5.3 17.6 22.1 37.5 17.88 50,000 0
@@
 ECE 11.7 11.2 11.1 12.6 12.3 7.7 5.4 6.5 8.8 3.4 2.9 21.9 3.5 3.7 4.0 8.46
-MEMO* Acc. 6.8 8.5 7.5 20.5 13.4 19.8 25.8 22.1 27.7 27.6 60.9 11.3 24.4 32.2 37.9 23.09 50,000$\times$65 50,000$\times$64
+[…]
 ECE 24.1 24.2 22.9 5.3 19.3 14.8 23.4 30.4 18.7 24.6 7.2 14.9 29.4 19.3 13.6 19.47
@@
 ECE 3.7 4.0 3.3 7.4 6.9 3.2 1.7 1.8 2.9 1.3 1.6 8.7 1.3 1.3 1.3 3.37
-EATA (Ours) Acc. 35.6 37.8 36.1 35.3 35.4 47.6 53.1 51.4 45.9 60.1 67.6 45.8 58.2 60.3 55.4 48.36
-ECE 10.5 10.2 10.3 12.9 12.3 8.3 6.6 6.7 8.2 4.7 4.5 9.9 5.1 5.3 5.6 8.08 50,000 26,188
-EATA-C (Ours) Acc. 37.2 39.2 38.0 35.5 35.9 48.2 52.3 51.9 46.5 60.4 67.3 48.7 58.0 60.5 56.0 49.04
-ECE 7.1 6.8 6.8 9.3 8.9 4.3 3.6 3.0 4.5 1.8 1.9 4.5 2.3 2.4 2.4 4.63 82,492 32,492
+Acc. 35.6 37.8 36.1 35.3 35.4 47.6 53.1 51.4 45.9 60.1 67.6 45.8 58.2 60.3 55.4 48.36
+EATA (Ours) ECE 10.5 10.2 10.3 12.9 12.3 8.3 6.6 6.7 8.2 4.7 4.5 9.9 5.1 5.3 5.6 8.08 50,000 26,188
+Acc. 37.2 39.2 38.0 35.5 35.9 48.2 52.3 51.9 46.5 60.4 67.3 48.7 58.0 60.5 56.0 49.04
+EATA-C (Ours) ECE 7.1 6.8 6.8 9.3 8.9 4.3 3.6 3.0 4.5 1.8 1.9 4.5 2.3 2.4 2.4 4.63 82,492 32,492
 ViT (LN) Source Acc. 12.9 17.6 11.7 34.4 27.7 43.7 36.2 43.4 45.4 52.8 73.3 45.5 37.9 54.7 60.2 39.84 50,000 0
@@
 ECE 24.1 10.1 11.7 8.6 10.4 7.5 12.4 8.0 8.0 4.8 2.4 5.2 10.0 4.3 4.1 8.78
-MEMO* Acc. 32.2 35.1 32.6 37.5 28.5 43.3 40.1 45.2 47.0 53.9 73.3 53.3 39.6 59.5 62.8 45.60 50,000$\times$65 50,000$\times$64
+[…]
 ECE 33.0 32.5 33.2 36.7 45.7 41.0 47.2 39.8 38.8 33.7 20.0 26.7 48.5 30.4 27.2 35.63
-CoTTA Acc. 45.6 46.8 28.1 45.3 42.0 57.6 49.8 59.5 57.6 65.2 74.9 58.2 55.5 66.3 65.7 54.54 50,000$\times$3 50,000
+[…]
 ECE 9.0 8.9 9.9 7.9 12.9 8.1 12.1 7.1 9.0 6.4 4.3 6.6 10.7 6.7 6.6 8.42
@@
 ECE 8.0 7.0 7.6 6.1 6.7 5.9 9.2 5.5 5.9 3.6 2.1 4.0 7.2 3.6 3.4 5.71
-EATA (Ours) Acc. 50.5 51.3 51.3 55.9 56.0 60.7 58.0 64.2 62.2 70.1 77.4 66.7 63.9 70.5 68.9 61.84
-ECE 10.6 10.1 10.3 9.4 8.8 7.8 9.1 6.4 7.3 5.4 2.8 6.1 6.6 4.5 4.8 7.34 50,000 32,524
-EATA-C (Ours) Acc. 56.6 57.6 57.2 57.8 59.8 64.6 64.4 69.1 67.2 73.3 78.6 67.0 70.2 73.9 71.2 65.90
-ECE 5.3 4.8 5.2 5.3 4.1 3.5 4.0 2.8 2.7 2.1 1.4 2.9 2.2 1.7 2.2 3.34 82,364 32,364
-TABLE E: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC in lifelong test-time adaptation scenario.
+Acc. 50.5 51.3 51.3 55.9 56.0 60.7 58.0 64.2 62.2 70.1 77.4 66.7 63.9 70.5 68.9 61.84
+EATA (Ours) ECE 10.6 10.1 10.3 9.4 8.8 7.8 9.1 6.4 7.3 5.4 2.8 6.1 6.6 4.5 4.8 7.34 50,000 32,524
+Acc. 56.6 57.6 57.2 57.8 59.8 64.6 64.4 69.1 67.2 73.3 78.6 67.0 70.2 73.9 71.2 65.90
+EATA-C (Ours) ECE 5.3 4.8 5.2 5.3 4.1 3.5 4.0 2.8 2.7 2.1 1.4 2.9 2.2 1.7 2.2 3.34 82,364 32,364
+TABLE B: Semantic segmentation results (mIoU in %) on the Cityscapes-to-ACDC in lifelong TTA scenario.
 The model is continually adapted to the four adverse conditions for ten rounds without resetting model parameters, i.e., lifelong adaptation.
@@
 CoTTA 70.9 41.1 62.4 59.7 70.9 41.0 62.5 59.7 70.9 40.8 62.6 59.7 70.8 40.6 62.6 59.7 70.8 40.6 62.6 59.7 cont.
+DAT 71.7 44.6 63.8 62.2 70.3 44.3 62.5 61.3 69.1 43.4 61.4 60.3 68.0 42.0 60.9 59.4 66.8 40.9 60.4 59.0 cont.
 EATA 69.1 40.5 59.8 58.1 69.3 41.1 60.0 58.4 69.3 41.5 60.1 58.6 69.3 41.8 60.1 58.6 69.2 42.1 59.9 58.5 cont.
@@
 CoTTA 65.2 34.3 56.9 52.4 64.6 33.4 55.9 51.6 63.9 32.4 54.7 50.6 63.2 31.5 53.7 49.6 52.7 30.4 52.6 48.7 58.4
+DAT 66.4 40.7 59.7 58.3 66.1 40.6 59.8 57.8 65.6 40.3 59.3 56.8 65.1 39.7 58.7 56.0 63.8 39.6 58.2 55.4 57.0
 EATA 69.0 42.3 59.7 58.2 68.8 42.5 59.4 57.9 68.6 42.7 58.9 57.4 68.3 42.8 58.4 56.9 67.9 42.8 57.7 56.3 57.0
 EATA-C 72.7 47.2 64.5 63.9 71.8 48.2 64.2 64.2 71.5 48.1 64.6 64.6 71.8 48.4 64.5 64.4 72.0 48.7 64.3 64.1 61.6
-TABLE F: Comparison on CIFAR-10-C.
-Results are evaluated in the single-domain adaptation scenario, and averaged over 15 different corruption types and 5 severity levels (totally 75).
-We use * to denote episodic adaptation.
-Model Acc. (%) #Forwards #Backwards
-ResNet-26 (GroupNorm) 77.5 10,000 0
-[…]
-[…]
-ResNet-26 (GroupNorm)+JT 77.2 10,000 0
-[…]
-[…]
-ResNet-26 (BatchNorm) 71.6 10,000 0
+TABLE C: Effects of components in EATA-C.
+Results obtained in single-domain TTA scenario, i.e., the model parameters are reset before adapting to a new corruption.
+CL denotes consistency loss.
+ER denotes entropy regularization.
+FR denotes Fisher regularization.
+$S$ denotes active sample selection.
+Experiment CL ER FR $S$ ResNet-50 over ImageNet-R
+Acc. ECE #Forwards #Backwards
+Source 38.0 17.7 30,000 0
 […]
@@
 […]
-Figure B: Comparison between ETA and Tent in terms of the number of backward propagation on ImageNet-C with different corruption types and severity levels.
-Figure C: Comparison of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50.
+[…]
+[…]
+[…]
+7 46.9 14.4 35,147 5,147
+8 44.6 11.9 34,991 4,991
+9 (ETA-C) 46.8 13.4 35,137 5,137
+10 47.1 13.7 35,179 5,179
+11 44.6 12.1 34,935 4,935
+12 (EATA-C) 47.1 13.3 35,122 5,122
+Experiment CL ER FR $S$ ViT-Base over ImageNet-C (level 5)
+Acc. ECE #Forwards #Backwards
+Source 39.8 7.5 50,000 0
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+7 65.6 4.4 83,182 33,182
+8 52.9 3.4 73,323 23,323
+9 (ETA-C) 65.3 3.4 82,015 32,015
+10 66.2 4.3 83,429 33,429
+11 53.2 3.4 73,723 23,723
+12 (EATA-C) 65.9 3.3 82,364 32,364
+TABLE D: Comparison of parameter update strategies.
+EATA-C (Ours) updates only the normalization layer parameters, while EATA-C updates all model parameters.
+Results are obtained on ImageNet-C (Gaussian, level 5) with ViT-Base.
+Loss denotes the average KL-Divergence between the full network and the sub-network across the dataset during online TTA.
+Time and peak memory usage are measured by processing 50,000 images.
+Method
+Loss
+Acc. (%)
+ECE (%)
+Time (s)
+Memory (MB)
+Source 0.565 12.9 14.2 55.7 816.6
+SAR - 43.1 8.0 167.3 7433.2
+EATA-C 0.437 56.0 5.9 140.9 9920.5
+EATA-C (Ours) 0.451 56.8 5.2 114.9 5786.6
+Figure D: Comparison between ETA and Tent w.r.t. the number of backward passes on ImageNet-C with different corruption types and severity levels.
+Figure E: Visualization of selected (reliable) and removed (unreliable) test samples for adaptation according to Eqn.
+(3).
+Predicition confidence is the highest class probability.
+Results are obtained on ImageNet-C(Gaussian, level 5) using the source ViT-Base model.
+Figure F: Comparison of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50.
 We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation).
 The model performs lifelong adaptation, in which the model parameters will never be reset.
-Figure D: Comparisons of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50.
+Figure G: Comparisons of preventing forgetting on ImageNet-C (severity levels 1-4) with ResNet-50.
 We record the OOD corruption accuracy on each corrupted test set and the associated ID clean accuracy (after OOD adaptation).
-The model parameters of both Tent and our EATA are reset before adapting to a new corruption type.
+The model performs single-domain adaptation, where model parameters of both Tent and our EATA are reset before adapting to a new corruption type.